### v1

# Accurate and thermodynamically consistent hydrogen equation of state for planetary modeling with flow matching

Abstract
Accurate determination of the equation of state of dense hydrogen is essential for understanding gas giants. Currently, there is still no consensus on methods for calculating its entropy, which play a fundamental role and can result in qualitatively different predictions for Jupiter’s interior. Here, we investigate various aspects of entropy calculation for dense hydrogen based on ab initio molecular dynamics simulations. Specifically, we employ the recently developed flow matching method to validate the accuracy of the traditional thermodynamic integration approach. We then clearly identify pitfalls in previous attempts and propose a reliable framework for constructing the hydrogen equation of state, which is accurate and thermodynamically consistent across a wide range of temperature and pressure conditions. This allows us to conclusively address the long-standing discrepancies in Jupiter’s adiabat among earlier studies, demonstrating the potential of our approach for providing reliable equations of state of diverse materials.
Introduction. The equation of state (EOS) of dense hydrogen (H) is one of the core ingredients to model the internal structure of giant gaseous planets, such as Jupiter and Saturn [1]. With the wealth of data from recent spacecraft missions, Galileo [2] and Juno [3], which measured Jupiter’s atmospheric composition and gravitational field, lack of accurate knowledge of the thermodynamic properties of hydrogen is now considered as one of the major roadblocks in planetary science [4, 5, 6, 7]. As parameterized planetary models already require several hypotheses [5, 8, 9, 10], an uncontested EOS is needed to advance the field.
Since the high temperature and pressure inside these planets are still largely beyond the reach of current experimental techniques, one usually relies instead on numerical simulations, typically ab initio molecular dynamics (MD) based on electron structure methods such as density functional theory (DFT) or quantum Monte Carlo (QMC). Up to now, there have been several widely used ab initio EOSs for planetary science applications, including the one by Chabrier, Mazevet and Soubiran (CMS19) [11], by Militzer and Hubbard (MH13) [12], and the Rostock EOS (REOS2 [13], REOS3 [14]) of various versions. All of them are derived from DFT-MD simulations using the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) [15]. Despite this common origin, however, they can make significant differences in the resulting Jupiter model, such as the location of adiabat and the predicted size of core mass [4, 7].
The key factor behind these discrepancies of planetary models is the calculation of entropy, which has been recognized and discussed to some extent in several studies [12, 4, 6]. Unlike observables such as energy and pressure , the Helmholtz free energy and entropy are not directly accessible from ab initio simulations and often rely on extra procedures called thermodynamic integration (TI). One common implementation of this technique involves interpolating the available data to create a continuous path on the temperature-density phase diagram [13, 4]. From basic statistical mechanics we have
These two equations correspond to integrating the energy (pressure) of the system along isochores (isotherms), respectively. However, this approach has been criticized because, due to limited number of simulation points in practice, the interpolation procedure may introduce uncontrollable systematic errors [12].
An alternative TI scheme is called the coupling constant integration (CCI) [16, 17] or Hamiltonian thermodynamic integration [18]. By connecting the target state to an artificial Hamiltonian with tractable thermodynamic properties, this method further allows for the calculation of absolute entropy. However, it is also much more expensive due to the need to perform additional MD simulations, and therefore only suitable for investigating a small region of the phase diagram [19, 12]. Besides, one has to carefully choose relevant parameters in the artificial system and MD simulations to ensure well-converged results [20], and the errors associated with discrete simulation points still remain.
The uncertainties involved in entropy calculation are further exacerbated by the fact that the EOS used for planetary modeling is usually composed of several parts, each based on different theoretical approach and has different phase region of validity. For example, although DFT-MD can provide a good description of electron correlations at intermediate temperatures and densities, especially near pressure-induced dissociation, it will quickly become inefficient at lower densities . For such a region deep within the molecular phase, empirical chemical models such as the well known Saumon–Chabrier–van Horn (SCvH) EOS [21] are generally believed to be accurate and reliable. In practice, the data from these different methods needs to be connected in some way. This, however, could possibly introduce critical errors, as we will demonstrate below.
In this Letter, we focus on the crucial issue of entropy calculation at the DFT-PBE level of theory, aiming to resolve long-standing disagreements among various Jupiter model predictions, which are highly sensitive to small changes in entropy [12, 4, 22, 6]. We rely solely on ab initio theory and well-established experimental data without invoking indirect observations. The purpose of this paper is not to reconcile planetary models with spacecraft measurements [1, 5], but rather to provide a non-parametric hydrogen EOS, ranging from roughly 1 bar to 700GPa, that forms a solid foundation for future model revisions.
We calculate the free energy and entropy of dense hydrogen using a recently developed method called flow matching [23, 24, 25]. Unlike the demonstrative example shown in Ref. [26], here the calculations are based on DFT-MD simulations and thus have ab initio accuracy. Basically, we employ a machine-learned invertible transformation , which acts on configuration samples of the proton coordinates, to directly connect two phase points of interest. Given the transformation can bring the two states close enough to ensure a sufficient overlap, their free energy difference (scaled by temperature) can then be efficiently estimated as follows:
where are the corresponding Boltzmann distributions, , being the Boltzmann constant, and
are known as the forward and reverse work, respectively [27]. In practice, the two estimators in Eq. (2) provide both upper and lower bound of the free energy difference, which can be pretty tight [27]. In this way, one can accurately estimate the entropy difference based solely on samples and energy functions of the two end states, thus avoid errors arising from any kinds of interpolation between them as done in TI. Moreover, it is also fairly straightforward to compute the absolute entropy by just choosing one of the state to have tractable thermodynamic properties.
The computational framework of Eqs. (2) and (3) is usually known as targeted free energy perturbation (TFEP) [28, 29]. Thanks to the rapid development of deep generative models in recent years, the desired invertible map can now be efficiently represented by a class of generative models called normalizing flow [30], making TFEP emerge as a promising tool for free energy calculation [31, 32, 33, 34]. In contrast to the “variational” approach, which has previously been applied to a broad range of many-body problems [35, 36, 32, 33, 37, 38, 39], the flow matching method we employ in this work represents a different training scheme of the flow model [23, 24, 25], which offers a variety of advantages in both accuracy and efficiency; see Supplemental Material [27] for details.
Entropy from ab initio data. We employ the flow matching method to study the ab initio portion of the phase diagram, with K and . Specifically, we repeatedly train a transformation and use Eq. (2) to compute the entropy difference for each neighboring pair of states on a grid; see Supplemental Material [27] for an example. We also calculate the absolute entropy of a reference state within the same framework and then shift the entropy of other states rigidly. In practice, we choose the reference point at and [27], where the computed value of the absolute entropy is in excellent agreement with those reported by Morales et al. [19] based on coupled electron-ion Monte Carlo (CEIMC) simulations.
Figure 1 shows our results (solid line) for several isotherms of the specific entropy and pressure , together with two widely used hydrogen EOSs in past literature: CMS19 [11] (dotted line) and REOS3 [14] (dashed line). For convenience, we also include an error bar of absolute entropy at the reference point, and the CEIMC entropy of Morales et al. [19] (dash-dotted line) at relatively high densities. Note the REOS3 entropy is calculated by us using TI (see Eq. (1)), and, for sake of visualization, we have introduced a global entropy offset to make it match our data at the bottom-left phase point, i.e., and . We do not compare with the MH13 EOS [12] since it is derived directly for hydrogen-helium mixtures.
One striking feature of Fig. 1(a) is the non-monotonic behavior of entropy at low temperatures in both our results and REOS3, which is qualitatively different from CMS19. Moreover, CMS19 does not feature the typical slope change of the relation, as shown in Fig. 1(b), which is emanating from the critical point of the liquid-liquid phase transition between molecular and atomic phase [40, 41]. This indicates that although derived from DFT-PBE data in the phase region considered here [11, 42, 43, 19], the CMS19 EOS does not closely capture the pressure-induced molecular dissociation, which could have a notable impact on Jupiter models [27].
The observed behavior of entropy and pressure in our results, as described above, can be understood by virtue of the thermodynamic relation [21]: . In fact, by examining the pressure isochores as illustrated in the inset of Fig. 1(b), the region where the pressure decreases with temperature corresponds precisely to the region of increasing entropy in Fig. 1(a). Physically, they are just different manifestations of the molecular dissociation effects: the attractive force from broken molecular bonds can dominates the repulsion from kinetic motion, while the larger number of configurations involving only individual H atoms (compared to those with bound H2-like pairs) can compensate for the smaller occupied volume caused by increasing density.
Benchmarking thermodynamic integration. We can use the flow matching entropy to benchmark the TI approach, which simply amounts to integrating Eq. (1) using our pressure and energy data. We find that the TI entropies are virtually indistinguishable from the solid lines in Fig. 1(a). Besides, they do not depend on the choice of integration path and hence show very good thermodynamic consistency. Contrary to the previous belief [12], our observations indicate that TI performs quite well for a discrete grid with spacings in the order of K and . When combined with the absolute entropy calculation on only a single referent point, as done above, we can then achieve the same level of accuracy as the “full” CCI approach in Ref. [12], but with far less computational costs.
Note that compared to REOS3, we have added some additional simulation points at relatively low densities to improve the characterization of molecular dissociation. Nevertheless, TI is not fundamentally flawed even with the relatively sparse grid of REOS3 in the ab initio region, and the resulting entropy remains generally consistent with our data, as shown by the dashed lines in Fig. 1(a). However, TI becomes truly problematic when we further extend the integration path to touch other regions based on inconsistent theories. We can quantitatively illustrate such inconsistency by performing the line integral along each closed local square loop of the tabular data, which should in principle vanish everywhere. However, for REOS3 we found large deviation from this ideal behavior between the ab initio region and the chemical model at lower densities [44, 45], as shown in Figure 2. This implies that the resulting entropy will be ambiguously defined and strongly depend on the choice of integration path and initial reference point; see Supplemental Material [27] for more details. One can then identify an entropy error of approximately [27], which yields qualitatively large differences on planetary models derived in the past using this method [13, 4].
A thermodynamically consistent construction. We have shown that performing a single TI over an EOS containing various inconsistent theories can lead to significant and uncontrolled errors of the entropy. A better approach is to perform TI on each component (i.e., region where the same theory applies) of the EOS separately, and then interpolate the resulting free energy function [27]. In this way, one can derive the pressure and energy near the interpolation boundaries by taking partial derivatives, which are naturally consistent by construction. Another advantage is that one can easily incorporate accurate values of the absolute entropy into each separate region. This ensures that any residual interpolation errors that one can make at the boundaries remain localized and do not propagate elsewhere.
Figure 3(a-d) show our final EOS (solid lines), which consist of the ab initio DFT-MD part at , the SCvH EOS at , and an interpolation region between them. For comparison, we also include the entropy and pressure isotherms of Miguel et al. [4] (dashed lines), which are obtained by a single TI over the entire phase region of REOS3. Note the new interpolation procedure depends very sensitively on the entropy difference between SCvH and the ab initio data [27]. In doing this, we then surprisingly found the reported (absolute) entropies of SCvH at room conditions (e.g., at standard pressure 1 bar and temperature 298.15K) are about (or ) higher than both the experimental data [46, 47] and theoretical calculations [48]. Only after such errors are fixed (by applying a global entropy shift on the SCvH side) can we achieve a high-quality interpolation that respects fundamental thermodynamic constraints; see Supplemental Material [27] for details. This observation, which results from totally independent experimental or numerical benchmarks at both low ( 1 bar) and high ( 700GPa) pressures, is a very strong evidence on the accuracy of our construction. Fig. 3(e) also shows the local loop integral values of our final EOS; compared to Fig. 2, one clearly sees the new interpolation approach can indeed yield much better thermodynamic consistency over the entire phase diagram, as discussed above.
Jupiter’s adiabat. We demonstrate the utility of the newly constructed hydrogen EOS by performing a preliminary calculation of Jupiter’s adiabat. We assume a homogeneous model with no compact core, and set a helium mass fraction of to ensure a fair comparison with previous works [4, 22, 12]. We adopt a simple linear mixing approximation for the entropy of mixtures, using the REOS3 EOS of helium derived by Miguel et al. in Ref. [22]. The starting point is bar and K [1], which corresponds to an entropy (for hydrogen) of .
Figure 4 shows our result for the adiabat (red dashed line), together with those derived from the REOS3 (black line) [22] and MH13 EOS (blue line) [12], respectively. For a comparison also with CMS19, see the Supplemental Material [27]. Note the hydrogen EOS used in the REOS3 adiabat was produced by Miguel et al. [4, 22] using a single TI over the entire phase region, as already shown by the dashed lines in Fig. 3. Our result clearly agrees well with the REOS3 adiabat up to about 1GPa, which is as expected since both the underlying hydrogen EOSs are derived at low pressures from the same SCvH model. However, at higher pressures, our adiabat begins to diverge from the REOS3 line and becomes cooler. This is a direct consequence of the systematically lower REOS3 entropy in the ab initio region in Fig. 3(a), which is in turn caused by the thermodynamic inconsistency shown in Fig. 2 [27]. Notice such a cooler adiabat passes very close to the miscibility boundary for hydrogen-helium mixtures, shown as the yellow line in Fig. 4, albeit derived from DFT-MD simulations using an exchange-correlation functional different from PBE [49]. This suggests an increased likelihood of phase separation between hydrogen and helium in Jupiter’s deep interior, causing the formation of so-called helium rain [1, 50, 51].
On the other hand, our adiabat at high pressures is apparently closer to the MH13 prediction, as shown by the blue line in Fig. 4. Note that the MH13 EOS is derived directly from ab initio MD simulations of a hydrogen-helium mixture and does not rely on the linear mixing approximation. Overall, we show that, at the DFT-PBE level of theory, the long-standing discrepancy between the Jupiter models of REOS3 and MH13 primarily stems from a problematic TI calculation of the REOS3 entropy that is not everywhere thermodynamically consistent. However, according to our earlier benchmark using flow matching, this definitely does not imply that TI is inherently inferior to the CCI approach for a finite number of simulation points, as speculated in Ref. [12]. Another observation from Fig. 4 is that the MH13 adiabat [12] was derived only for the ab initio region at high pressures. In fact, Ref. [12] did not really attempt a thermodynamic consistent interpolation with other regions to construct a complete EOS, which, however, is definitely needed for a thorough study of planetary models.
Summary and outlook. In this Letter, we closely examine the entropy calculation of dense hydrogen by benchmarking TI with an independent flow matching method. Based on this, we then construct a new hydrogen EOS that is both accurate and thermodynamically consistent over several orders of magnitude of pressure, which is crucial for reliable and conclusive planetary model predictions. Notably, the interpolation quality at the boundaries serves as a highly sensitive and stringent test for the accuracy of the theories on both sides [27]. Indeed, the new approach has allowed us to identify and correct a significant discrepancy in SCvH relative to experimental data at low pressures, which is fairly surprising and remarkable given the chemical model’s long history of nearly 30 years.
We conclusively address the long-standing issue of entropy disagreements among various popular EOSs used in planetary modeling [11, 12, 14], all of which are derived from the same DFT-PBE level of electronic structure theory. To further improve the hydrogen EOS and potentially reconcile models with observational data [5], one can upgrade the description of electronic correlation from DFT-PBE to a higher-level theory [52, 53, 54, 39]. We also note that as a generic tool, flow matching can also be used to study other rich phases of hydrogen [55, 56] or other interesting condensed matter systems [51, 57], where accurate calculation of free energy plays a significant role.
Acknowledgements.
We acknowledge useful discussions with Lei Wang, Ravit Helled, Gilles Chabrier, and Armin Bergermann. We also acknowledge Cesare Cozza for providing initial data and support with the DFT calculations. H.X. and G.M. acknowledge financial support from the Swiss National Science Foundation (grant PCEFP2_203455). S.H. acknowledges financial support from the Swiss National Science Foundation (grant 200020_215634).References
- Helled et al. [2020] R. Helled, G. Mazzola, and R. Redmer, Understanding dense hydrogen at planetary conditions, Nature Reviews Physics 2, 562 (2020).
- Johnson et al. [1992] T. V. Johnson, C. M. Yeates, and R. Young, Space science reviews volume on Galileo Mission overview, Space Science Reviews 60, 3 (1992).
- Bolton et al. [2017] S. J. Bolton, J. Lunine, D. Stevenson, J. E. P. Connerney, S. Levin, T. C. Owen, F. Bagenal, D. Gautier, A. P. Ingersoll, G. S. Orton, T. Guillot, W. Hubbard, J. Bloxham, A. Coradini, S. K. Stephens, P. Mokashi, R. Thorne, and R. Thorpe, The Juno Mission, Space Science Reviews 213, 5 (2017).
- Miguel, Y. et al. [2016] Miguel, Y., Guillot, T., and Fayon, L., Jupiter internal structure: the effect of different equations of state, A&A 596, A114 (2016).
- Debras and Chabrier [2019] F. Debras and G. Chabrier, New Models of Jupiter in the Context of Juno and Galileo, The Astrophysical Journal 872, 100 (2019).
- Mazevet, S. et al. [2022] Mazevet, S., Licari, A., and Soubiran, F., Benchmarking the ab initio hydrogen equation of state for the interior structure of jupiter, A&A 664, A112 (2022).
- Howard, S. et al. [2023] Howard, S., Guillot, T., Bazot, M., Miguel, Y., Stevenson, D. J., Galanti, E., Kaspi, Y., Hubbard, W. B., Militzer, B., Helled, R., Nettelmann, N., Idini, B., and Bolton, S., Jupiter’s interior from Juno: Equation-of-state uncertainties and dilute core extent, A&A 672, A33 (2023).
- Nettelmann et al. [2021] N. Nettelmann, N. Movshovitz, D. Ni, J. J. Fortney, E. Galanti, Y. Kaspi, R. Helled, C. R. Mankovich, and S. Bolton, Theory of Figures to the Seventh Order and the Interiors of Jupiter and Saturn, The Planetary Science Journal 2, 241 (2021).
- Miguel, Y. et al. [2022] Miguel, Y., Bazot, M., Guillot, T., Howard, S., Galanti, E., Kaspi, Y., Hubbard, W. B., Militzer, B., Helled, R., Atreya, S. K., Connerney, J. E. P., Durante, D., Kulowski, L., Lunine, J. I., Stevenson, D., and Bolton, S., Jupiter’s inhomogeneous envelope, A&A 662, A18 (2022).
- Militzer et al. [2022] B. Militzer, W. B. Hubbard, S. Wahl, J. I. Lunine, E. Galanti, Y. Kaspi, Y. Miguel, T. Guillot, K. M. Moore, M. Parisi, J. E. P. Connerney, R. Helled, H. Cao, C. Mankovich, D. J. Stevenson, R. S. Park, M. Wong, S. K. Atreya, J. Anderson, and S. J. Bolton, Juno Spacecraft Measurements of Jupiter’s Gravity Imply a Dilute Core, The Planetary Science Journal 3, 185 (2022).
- Chabrier et al. [2019] G. Chabrier, S. Mazevet, and F. Soubiran, A new equation of state for dense hydrogen–helium mixtures, The Astrophysical Journal 872, 51 (2019).
- Militzer and Hubbard [2013] B. Militzer and W. B. Hubbard, Ab initio equation of state for hydrogen–helium mixtures with recalibration of the giant-planet mass–radius relation, The Astrophysical Journal 774, 148 (2013).
- Nettelmann et al. [2012] N. Nettelmann, A. Becker, B. Holst, and R. Redmer, Jupiter models with improved ab initio hydrogen equation of state (h-reos.2), The Astrophysical Journal 750, 52 (2012).
- Becker et al. [2014] A. Becker, W. Lorenzen, J. J. Fortney, N. Nettelmann, M. Schöttler, and R. Redmer, Ab initio equations of state for hydrogen (h-reos.3) and helium (he-reos.3) and their implications for the interior of brown dwarfs, The Astrophysical Journal Supplement Series 215, 21 (2014).
- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
- Kirkwood [1935] J. G. Kirkwood, Statistical Mechanics of Fluid Mixtures, The Journal of Chemical Physics 3, 300 (1935).
- de Wijs et al. [1998] G. A. de Wijs, G. Kresse, and M. J. Gillan, First-order phase transitions by first-principles free-energy calculations: The melting of al, Phys. Rev. B 57, 8223 (1998).
- Frenkel and Smit [2023] D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications, 3rd ed. (Academic Press, 2023).
- Morales et al. [2010a] M. A. Morales, C. Pierleoni, and D. M. Ceperley, Equation of state of metallic hydrogen from coupled electron-ion monte carlo simulations, Phys. Rev. E 81, 021202 (2010a).
- Bergermann et al. [2024] A. Bergermann, M. French, and R. Redmer, Ab initio calculation of the miscibility diagram for mixtures of hydrogen and water, Phys. Rev. B 109, 174107 (2024).
- Saumon et al. [1995] D. Saumon, G. Chabrier, and H. M. van Horn, An Equation of State for Low-Mass Stars and Giant Planets, Astrophysical Journal Supplement 99, 713 (1995).
- Miguel, Y. et al. [2018] Miguel, Y., Guillot, T., and Fayon, L., Jupiter internal structure: the effect of different equations of state (corrigendum), A&A 618, C2 (2018).
- Albergo and Vanden-Eijnden [2023] M. S. Albergo and E. Vanden-Eijnden, Building normalizing flows with stochastic interpolants, in The Eleventh International Conference on Learning Representations (2023).
- Lipman et al. [2023] Y. Lipman, R. T. Q. Chen, H. Ben-Hamu, M. Nickel, and M. Le, Flow matching for generative modeling, in The Eleventh International Conference on Learning Representations (2023).
- Liu et al. [2023] X. Liu, C. Gong, and qiang liu, Flow straight and fast: Learning to generate and transfer data with rectified flow, in The Eleventh International Conference on Learning Representations (2023).
- Zhao and Wang [2023] L. Zhao and L. Wang, Bounding free energy difference with flow matching, Chinese Physics Letters 40, 120201 (2023).
- [27] See Supplemental Material for more details on the theoretical formulations of targeted free energy perturbation and flow matching, as well as their implementation and some illustrative examples in practice. There are also further discussions about the effects of thermodynamic inconsistency on entropy calculation, our new EOS construction approach, and the comparison between Jupiter’s adiabats derived from various hydrogen EOSs.
- Jarzynski [2002] C. Jarzynski, Targeted free energy perturbation, Phys. Rev. E 65, 046122 (2002).
- Hahn and Then [2009] A. M. Hahn and H. Then, Using bijective maps to improve free-energy estimates, Phys. Rev. E 79, 011113 (2009).
- Papamakarios et al. [2021] G. Papamakarios, E. Nalisnick, D. J. Rezende, S. Mohamed, and B. Lakshminarayanan, Normalizing flows for probabilistic modeling and inference, Journal of Machine Learning Research 22, 1 (2021).
- Wirnsberger et al. [2020] P. Wirnsberger, A. J. Ballard, G. Papamakarios, S. Abercrombie, S. Racanière, A. Pritzel, D. Jimenez Rezende, and C. Blundell, Targeted free energy estimation via learned mappings, The Journal of Chemical Physics 153, 144112 (2020).
- Wirnsberger et al. [2022] P. Wirnsberger, G. Papamakarios, B. Ibarz, S. Racanière, A. J. Ballard, A. Pritzel, and C. Blundell, Normalizing flows for atomic solids, Machine Learning: Science and Technology 3, 025009 (2022).
- Ahmad and Cai [2022] R. Ahmad and W. Cai, Free energy calculation of crystalline solids using normalizing flows, Modelling and Simulation in Materials Science and Engineering 30, 065007 (2022).
- Molina-Taborda et al. [2024] A. Molina-Taborda, P. Cossio, O. Lopez-Acevedo, and M. Gabrié, Active Learning of Boltzmann Samplers and Potential Energies with Quantum Mechanical Accuracy, Journal of Chemical Theory and Computation 20, 8833 (2024).
- Li and Wang [2018] S.-H. Li and L. Wang, Neural network renormalization group, Phys. Rev. Lett. 121, 260601 (2018).
- Wu et al. [2019] D. Wu, L. Wang, and P. Zhang, Solving statistical mechanics using variational autoregressive networks, Phys. Rev. Lett. 122, 080602 (2019).
- Xie et al. [2022] H. Xie, L. Zhang, and L. Wang, Ab-initio study of interacting fermions at finite temperature with neural canonical transformation, Journal of Machine Learning 1, 38 (2022).
- Xie et al. [2023a] H. Xie, L. Zhang, and L. Wang, of two-dimensional electron gas: A neural canonical transformation study, SciPost Phys. 14, 154 (2023a).
- Xie et al. [2023b] H. Xie, Z.-H. Li, H. Wang, L. Zhang, and L. Wang, Deep variational free energy approach to dense hydrogen, Phys. Rev. Lett. 131, 126501 (2023b).
- Lorenzen et al. [2010] W. Lorenzen, B. Holst, and R. Redmer, First-order liquid-liquid phase transition in dense hydrogen, Phys. Rev. B 82, 195107 (2010).
- Geng et al. [2019] H. Y. Geng, Q. Wu, M. Marqués, and G. J. Ackland, Thermodynamic anomalies and three distinct liquid-liquid transitions in warm dense liquid hydrogen, Phys. Rev. B 100, 134109 (2019).
- Caillabet et al. [2011] L. Caillabet, S. Mazevet, and P. Loubeyre, Multiphase equation of state of hydrogen from ab initio calculations in the range 0.2 to 5 g/cc up to 10 ev, Phys. Rev. B 83, 094101 (2011).
- Holst et al. [2008] B. Holst, R. Redmer, and M. P. Desjarlais, Thermophysical properties of warm dense hydrogen using quantum molecular dynamics simulations, Phys. Rev. B 77, 184201 (2008).
- Juranek et al. [2002] H. Juranek, R. Redmer, and Y. Rosenfeld, Fluid variational theory for pressure dissociation in dense hydrogen: Multicomponent reference system and nonadditivity effects, The Journal of Chemical Physics 117, 1768 (2002).
- Holst et al. [2007] B. Holst, N. Nettelmann, and R. Redmer, Equation of state for dense hydrogen and plasma phase transition, Contributions to Plasma Physics 47, 368 (2007).
- Cox et al. [1989] J. D. Cox, D. D. Wagman, and V. A. Medvedev, Codata key values for thermodynamics, in Series on Thermodynamic Properties (1989).
- Lemmon et al. [2023] E. W. Lemmon, I. H. Bell, M. L. Huber, and M. O. McLinden, Thermophysical properties of fluid systems, in NIST Chemistry WebBook, NIST Standard Reference Database Number 69, edited by P. Linstrom and W. Mallard (National Institute of Standards and Technology, Gaithersburg MD, 20899, 2023).
- NIS [2022] NIST Computational Chemistry Comparison and Benchmark Database, NIST Standard Reference Database Number 101 (2022).
- Schöttler and Redmer [2018] M. Schöttler and R. Redmer, Ab initio calculation of the miscibility diagram for hydrogen-helium mixtures, Phys. Rev. Lett. 120, 115703 (2018).
- Vorberger et al. [2007] J. Vorberger, I. Tamblyn, B. Militzer, and S. A. Bonev, Hydrogen-helium mixtures in the interiors of giant planets, Phys. Rev. B 75, 024206 (2007).
- Chang et al. [2024] X. Chang, B. Chen, Q. Zeng, H. Wang, K. Chen, Q. Tong, X. Yu, D. Kang, S. Zhang, F. Guo, Y. Hou, Z. Zhao, Y. Yao, Y. Ma, and J. Dai, Theoretical evidence of h-he demixing under jupiter and saturn conditions, Nature Communications 15, 8543 (2024).
- Pierleoni et al. [2016] C. Pierleoni, M. A. Morales, G. Rillo, M. Holzmann, and D. M. Ceperley, Liquid–liquid phase transition in hydrogen by coupled electron–ion monte carlo simulations, Proceedings of the National Academy of Sciences 113, 4953 (2016).
- Mazzola et al. [2018] G. Mazzola, R. Helled, and S. Sorella, Phase diagram of hydrogen and a hydrogen-helium mixture at planetary conditions by quantum monte carlo simulations, Phys. Rev. Lett. 120, 025701 (2018).
- Nakano et al. [2020] K. Nakano, C. Attaccalite, M. Barborini, L. Capriotti, M. Casula, E. Coccia, M. Dagrada, C. Genovese, Y. Luo, G. Mazzola, A. Zen, and S. Sorella, TurboRVB: A many-body toolkit for ab initio electronic simulations by quantum Monte Carlo, The Journal of Chemical Physics 152, 204121 (2020).
- Morales et al. [2010b] M. A. Morales, C. Pierleoni, E. Schwegler, and D. M. Ceperley, Evidence for a first-order liquid-liquid transition in high-pressure hydrogen from ab initio simulations, Proceedings of the National Academy of Sciences 107, 12799 (2010b).
- Niu et al. [2023] H. Niu, Y. Yang, S. Jensen, M. Holzmann, C. Pierleoni, and D. M. Ceperley, Stable solid molecular hydrogen above 900 k from a machine-learned potential trained with diffusion quantum monte carlo, Phys. Rev. Lett. 130, 076102 (2023).
- Wang et al. [2023] X. Wang, Z. Wang, P. Gao, C. Zhang, J. Lv, H. Wang, H. Liu, Y. Wang, and Y. Ma, Data-driven prediction of complex crystal structures of dense lithium, Nature Communications 14, 2924 (2023).
- Zwanzig [1954] R. W. Zwanzig, High‐Temperature Equation of State by a Perturbation Method. I. Nonpolar Gases, The Journal of Chemical Physics 22, 1420 (1954).
- Bennett [1976] C. H. Bennett, Efficient estimation of free energy differences from monte carlo data, Journal of Computational Physics 22, 245 (1976).
- Chen et al. [2018] R. T. Q. Chen, Y. Rubanova, J. Bettencourt, and D. Duvenaud, Neural ordinary differential equations, in Proceedings of the 32nd International Conference on Neural Information Processing Systems, NIPS’18 (Curran Associates Inc., Red Hook, NY, USA, 2018) p. 6572–6583.
- Zhang et al. [2018] L. Zhang, W. E, and L. Wang, Monge-ampère flow for generative modeling (2018), arXiv:1809.10188 [cs.LG] .
- Vaswani et al. [2017] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, L. u. Kaiser, and I. Polosukhin, Attention is all you need, in Advances in Neural Information Processing Systems, Vol. 30, edited by I. Guyon, U. V. Luxburg, S. Bengio, H. Wallach, R. Fergus, S. Vishwanathan, and R. Garnett (Curran Associates, Inc., 2017).
- Köhler et al. [2019] J. Köhler, L. Klein, and F. Noé, Equivariant flows: sampling configurations for multi-body systems with symmetric energies (2019), arXiv:1910.00753 [stat.ML] .
- Satorras et al. [2021] V. G. Satorras, E. Hoogeboom, and M. Welling, E(n) equivariant graph neural networks, in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 139, edited by M. Meila and T. Zhang (PMLR, 2021) pp. 9323–9332.
- Taddei et al. [2015] M. Taddei, M. Ruggeri, S. Moroni, and M. Holzmann, Iterative backflow renormalization procedure for many-body ground-state wave functions of strongly interacting normal fermi liquids, Phys. Rev. B 91, 115106 (2015).
- Pfau et al. [2020] D. Pfau, J. S. Spencer, A. G. D. G. Matthews, and W. M. C. Foulkes, Ab initio solution of the many-electron schrödinger equation with deep neural networks, Phys. Rev. Res. 2, 033429 (2020).
- Pescia et al. [2022] G. Pescia, J. Han, A. Lovato, J. Lu, and G. Carleo, Neural-network quantum states for periodic systems in continuous space, Phys. Rev. Res. 4, 023138 (2022).
- Wilson et al. [2023] M. Wilson, S. Moroni, M. Holzmann, N. Gao, F. Wudarski, T. Vegge, and A. Bhowmik, Neural network ansatz for periodic wave functions and the homogeneous electron gas, Phys. Rev. B 107, 235139 (2023).
- Cassella et al. [2023] G. Cassella, H. Sutterud, S. Azadi, N. D. Drummond, D. Pfau, J. S. Spencer, and W. M. C. Foulkes, Discovering quantum phase transitions with fermionic neural networks, Phys. Rev. Lett. 130, 036401 (2023).
- Pescia et al. [2024] G. Pescia, J. Nys, J. Kim, A. Lovato, and G. Carleo, Message-passing neural quantum states for the homogeneous electron gas, Phys. Rev. B 110, 035108 (2024).
- Giannozzi et al. [2009] P. Giannozzi, S. Baroni, N. Bonini, M. Calandra, R. Car, C. Cavazzoni, D. Ceresoli, G. L. Chiarotti, M. Cococcioni, I. Dabo, A. Dal Corso, S. de Gironcoli, S. Fabris, G. Fratesi, R. Gebauer, U. Gerstmann, C. Gougoussis, A. Kokalj, M. Lazzeri, L. Martin-Samos, N. Marzari, F. Mauri, R. Mazzarello, S. Paolini, A. Pasquarello, L. Paulatto, C. Sbraccia, S. Scandolo, G. Sclauzero, A. P. Seitsonen, A. Smogunov, P. Umari, and R. M. Wentzcovitch, Quantum espresso: a modular and open-source software project for quantum simulations of materials, Journal of Physics: Condensed Matter 21, 395502 (19pp) (2009).
- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
- Monkhorst and Pack [1976] H. J. Monkhorst and J. D. Pack, Special points for brillouin-zone integrations, Phys. Rev. B 13, 5188 (1976).
- Bussi et al. [2007] G. Bussi, D. Donadio, and M. Parrinello, Canonical sampling through velocity rescaling, The Journal of Chemical Physics 126, 014101 (2007).
- Chen and Lipman [2024] R. T. Q. Chen and Y. Lipman, Flow matching on general geometries, in The Twelfth International Conference on Learning Representations (2024).
- Klein et al. [2023] L. Klein, A. Krämer, and F. Noe, Equivariant flow matching, in Thirty-seventh Conference on Neural Information Processing Systems (2023).
- Kuhn [1955] H. W. Kuhn, The hungarian method for the assignment problem, Naval Research Logistics Quarterly 2, 83 (1955).
- Militzer [2013] B. Militzer, Equation of state calculations of hydrogen-helium mixtures in solar and extrasolar giant planets, Phys. Rev. B 87, 014202 (2013).
- Becker et al. [2013] A. Becker, N. Nettelmann, B. Holst, and R. Redmer, Isentropic compression of hydrogen: Probing conditions deep in planetary interiors, Phys. Rev. B 88, 045122 (2013).
- Morales et al. [2013] M. A. Morales, J. M. McMahon, C. Pierleoni, and D. M. Ceperley, Nuclear quantum effects and nonlocal exchange-correlation functionals applied to liquid hydrogen at high pressure, Phys. Rev. Lett. 110, 065702 (2013).
- Chabrier and Potekhin [1998] G. Chabrier and A. Y. Potekhin, Equation of state of fully ionized electron-ion plasmas, Phys. Rev. E 58, 4941 (1998).
- Militzer and Ceperley [2000] B. Militzer and D. M. Ceperley, Path integral monte carlo calculation of the deuterium hugoniot, Phys. Rev. Lett. 85, 1890 (2000).
- Militzer and Ceperley [2001] B. Militzer and D. M. Ceperley, Path integral monte carlo simulation of the low-density hydrogen plasma, Phys. Rev. E 63, 066404 (2001).
Supplemental Material
Consider a hydrogen system composed of protons and electrons in a periodic box of volume . By introducing the dimensionless Wigner-Seitz parameter , where is the Bohr radius, one can then write the mass density (in unit of ) of the system as . Let be two state points on the temperature-density phase diagram, the corresponding Boltzmann distributions then read , respectively. denotes the proton coordinates, , and is the Boltzmann constant. is known as the partition function and directly related to the Helmholtz free energy .
In this work, we use the flow matching method to benchmark the calculation of free energy difference (scaled by temperature) between the two states, , against the conventional thermodynamic integration (TI) approach. Below we will discuss various aspect of the new method, including the underlying theoretical framework, training objectives, implementation details, and some illustrative examples of the results.
Appendix A Targeted free energy perturbation
In contrast to TI, flow matching is built on the framework of targeted free energy perturbation (TFEP) [28, 29], as already mentioned in the main text. Compared to the original free energy perturbation based on importance sampling [58], TFEP further leverages an invertible map in the configuration space to significantly increase state overlap and thereby reduces the variance of relevant estimators.
More specifically, consider a map that transforms coordinate samples of the state 0 into a new state , whose probability density function is given by the change-of-variables formula:
where the transformed “energy function” is defined as
Similarly, the inverse of would transform the state 1 to a new state with energy function
In practice, one would like to find an appropriate such that the transformed state has significant overlap with the target state . Based on the idea of importance sampling, the free energy difference can then be efficiently estimated as follows:
where
are known as the forward and reverse work, respectively. These results are precisely the Eqs. (2) and (3) in the main text.
Note that due to limited number of samples in practice, the two estimators in Eq. (S4) are actually biased, which can be readily demonstrated using Jensen’s inequality [29]. Specifically, the estimator involving the forward (reverse) work turns out to be an upper (lower) bound of the true free energy difference value. The extent of bias will, of course, depend on the quality of the map . In this work, we found these two bounds are generally quite tight; see the sections below for some illustrative examples. As a result, we choose to simply perform an arithmetic mean of the two bounds to generate the final data, although some more sophisticated and potentially superior approaches may also be considered [59]. Another observation is that the estimation of absolute entropy appears to be more straightforward within the TFEP framework compared to TI: once one of the two states is chosen to have tractable entropy, the practical implementation remains almost unchanged.
Appendix B Flow matching
As mentioned in the main text, one appealing feature of the TFEP estimators in Eq. (S4) is that they depend solely on samples from the two end states, as opposed to TI which requires information along an entire path. The price one pays, however, is the design of a suitable map , which is often challenging and lack of generic physical guidelines. This is arguably the main reason why the TFEP framework has not been widely used in realistic problems since its introduction.
The recent progress in the field of deep generative modeling has largely changed the situation. Specifically, the desired invertible map can be efficiently represented by a class of generative models called normalizing flow [30], making TFEP reemerge as a promising tool for free energy calculation. In previous attempts [31, 32, 33, 34], the training objectives have been exclusively based on variants of the Kullback-Leibler (KL) divergence, which provide direct access to the free energy difference as a variational upper or lower bound. This variational training scheme has been successfully applied to a broad range of physical problems, including lattice models [35, 36], atomic solids [32, 33], quantum dots [37], uniform electron gases [38], and also dense hydrogen [39]. However, it still has some drawbacks for the present purpose. First, the training process typically requires a large number of energy calculations of the underlying system, which would be time-consuming for ab initio methods like DFT. As a result, many works have used empirical [31, 32, 33] or machine-learned potentials [34], which would inevitably introduce new source of errors. Second, symmetry constraints often play an essential role in the efficient training of flow model, especially for systems like pure hydrogen. Although several neural network designs have been proposed to meet various symmetry requirements [31, 32], they are still not satisfactory enough regarding a proper balance between the expressive power and computational cost. The continuous normalizing flow (CNF) [60, 61] is arguably the most convenient and flexible candidate now to incorporate symmetries, but its variational training requires expensive numerical integrations of ordinary differential equation (ODE) and scales unfavorably for large system sizes.
The recently developed flow matching method [23, 24, 25] has largely overcome these difficulties. More specifically, one can smoothly connect the two end probability distributions by introducing an interpolant , where is a continuous “time” parameter, such as
The essence is that the velocity field underlying such an interpolated “probability path” can then be obtained by minimization of the following objective:
Notice this expression requires only samples from the two end distributions; there are not any on-the-fly evaluation of configurational energy or ODE integration during the training stage. On the other hand, these types of costly computations are indeed present when estimating the free energy difference after the training is finished, as shown in Eqs. (S4) and (S5). In particular, the transformed coordinates and corresponding log-Jacobian determinant should be evaluated by jointly integrating the following set of ODEs [60, 61]:
However, it should be noted that such computations are performed only once and thus still practically affordable.
As mentioned above, compared to previous network designs for discrete flow model [31, 32], the incorporation of various symmetries can also be more convenient for continuous normalizing flow. Specifically, the invariance of probability distribution can simply be achieved by requiring the underlying velocity field to be equivariant under the desired symmetry operation :
To do this in practice, we can leverage many recent advances in various areas, such as natural language processing [62], molecular simulation [63, 31], graph neural networks [64], and quantum many body computations [65, 66, 38, 67, 39, 68, 69, 70].
Appendix C Implementation details
The MD generation of coordinate samples used for training Eq. (S7) and the energy computation appearing in the free energy estimators Eq. (S4) are both performed in the DFT level using the QUANTUM ESPRESSO code [71]. We use the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) [15] and the pseudopotential of projector-augmented wave [72] type. We choose the system size , and the -space integration is performed on a Monkhorst-Pack grid [73]. The energy cutoff for the plane wave basis set and charge densities are and Ry, respectively. These parameters have been checked to yield well converged energetics. For MD simulations, the ion temperature is controlled using stochastic-velocity rescaling [74] and the time step is set to 12 a.u..
The network we use for the velocity field is adapted from FermiNet [66], which is permutation and translation equivariant by construction. We also modify the pair distance features to comply with the periodic nature of the simulation box [38, 67, 68, 69]. We choose a simple linear interpolating function
with the caveat in mind that we are working on a periodic box. This setting turns out to be equivalent to the Riemannian flow matching on a torus as proposed in Ref. [75]. To allow for more efficient use of the training data, we choose to further incorporate permutation symmetry explicitly into the flow matching objective Eq. (S7) itself [76]. In practice, this can be achieved by finding the permutation with minimal distance for each pair of samples using the Hungarian algorithm [77].
Appendix D Illustrative examples of the results
To showcase the utility of TFEP and flow matching in practice, we consider as a typical example the two end states with , and , , respectively. These states are near the molecular dissociation where the radial distribution function changes dramatically, as shown in Figure S1. Note we have measured the distance in terms of ; under such a scale, the simulation box size is and does not depend on the density of the system. This would facilitate the interpolation between the two Boltzmann distributions as done in the flow matching method. Fig. S1 also shows the radial distribution function of the new state transformed from state 0, obtained by integrating the first ODE in Eq. (S8) using the trained velocity field. The result is clearly very close to the target state 1, indicating a significant state overlap.
After the training is complete, one can further use Eq. (S4) to estimate the upper and lower bound of free energy difference between the two states, as shown by the red and black vertical line in Fig. S2, respectively. The distance between these two bounds is roughly within their own error bars, which implies our estimation have been tight enough. Figure S2 also shows histograms for the distribution of forward and reverse work , as defined in Eq. (S5). The significant overlap between these two work distributions indicates the high quality of the trained velocity field in effectively connecting the two end states being considered. Furthermore, one can clearly see the estimated free energy difference values coincide well with the intersection point of the two work distributions, where the probability densities are equal. This is a direct consequence of the fluctuation theorem as derived in Ref. [29].
We also use the same workflow outlined above to calculate the absolute free energy and entropy of the reference state at and , as mentioned in the main text. In practice, we choose the “tractable state” to correspond to a simple uniform distribution. The resulting radial distribution functions and free energy bounds are shown in Figures S3 and S4, respectively. Notice that in the present case, the two end states differ more substantially than those in Fig. S1, making the flow matching training of the velocity field more challenging. As a result, the forward and reverse work distributions also have less overlap, and the upper and lower bounds for the free energy difference are less tight compared to Fig. S2. Nonetheless, the absolute entropy of the reference state can still be estimated with a relative error of , which is comparable to the results reported by Morales et al. [19] using Hamiltonian thermodynamic integration and already satisfactory for the present purpose (see Fig. 1(a) in the main text).
Appendix E Effect of thermodynamic inconsistency on entropy calculation
In Fig. 2 of the main text, we have shown the entropy calculated by a single TI over the entire phase region of REOS3 can be thermodynamically inconsistent in the intermediate region between different theories. By comparing the maximum value of the color bar in Fig. 2 to the scale of Fig. 1(a) in the main text, one can conclude that a local inconsistency of such a magnitude will likely lead to significant discrepancy of the entropy. To illustrate this, we compute the REOS entropy by integrating either along the isotherm or isochore first, as shown by the red and blue solid lines in Figure S5, respectively. Note for a given grid of points, we have selected the one with lowest temperature and density as the reference point for integration. One can see for a large phase region (the main panel of Fig. S5) that covers both the chemical model and ab initio data, the results from the two integration paths begin to diverge significantly at intermediate densities, which is a direct consequence of the inconsistency observed in Fig. 2 of the main text. On the other hand, such a divergence becomes nearly negligible when the integration is performed exclusively over the ab initio region, as shown in the inset of Fig. S5.
The subtleties involved in the implementation of TI arise not only from different choices of integration path, but also the reference point. In fact, the latter can be seen as an alternative manifestation of the former, which can be understood by noting that an integration path can belong to different types (i.e., isotherm or isochore first) when viewed by different reference points. As an illustration, Fig. S5 also shows the entropy reported by Miguel et al. [4] (dashed lines). According to the authors, they processed the original REOS3 data by integrating along isotherm first, hence in accordance with the red solid lines of our result. However, these two sets of entropies clearly differ by an offset in the ab initio region of the main panel. The remedy of such an apparent disagreement turns out to be surprisingly easy: notice all the entropy isochores already align well, we just need to locate the reference point on another “correct” isotherm, say, 15000K. After doing this, the red solid lines then coincide perfectly with the dashed lines from Ref. [4], as shown in Figure S6.
Our reproduction of the entropy results by Miguel et al. [4] based on a specific integration path and reference point, as done above, demonstrates how fragile the “global” TI approach can be for an EOS with significant thermodynamic inconsistencies. Notice the solid and dashed lines in Fig. S5 differ by roughly , which can make a big difference on planetary model predictions as shown in the main text.
Appendix F More details on the thermodynamically consistent construction of our EOS
To construct an EOS that is thermodynamically consistent across several distinct theories, such as ab initio DFT and SCvH, one needs to first calculate the free energy and entropy of each region separately using the standard TI approach. We then perform a two-dimensional spline interpolation over each set of free energy data, following a similar approach to Militzer and Hubbard [12]. This in principle allows us to reproduce the original pressure and energy data by performing partial derivatives. The quality of such reproduction is, of course, closely related to the original level of thermodynamic consistency in each region. Once the free energy interpolation for each separate region is ready, we can then glue them together by performing another spline interpolation along their boundaries. Note one can utilize relevant partial derivative values derived from the free energy interpolations on both sides to ensure a smoother connection.
In this work, we attempted an interpolation between the ab initio DFT-PBE data at and the SCvH EOS at . Note we have chosen the locations of the density boundaries based on reasonable assumptions about the range of validity of both theories. Intuitively speaking, the size of the intermediate density gap reflects a trade-off between our confidence in available (but potentially conflicting) data and the flexibility to perform interpolations between them.
In principle, the free energy interpolation would be perfectly smooth if the absolute entropy and energy (measured relative to the same baseline) were exact on both sides. However, this is not the case in practice. First, it is generally believed that only the relative energies produced by DFT (using any given exchange-correlation functional) are reliable, rather than the absolute ones. This indicates that we are allowed to globally shift the DFT energies as we want, which, of course, leaves the entropy (and any other thermodynamic properties) unchanged. Second, we have surprisingly found the SCvH entropies are about (or ) higher than well-established experimental data at room conditions [46, 47], as mentioned in the main text. As shown in Figure S7, this discrepancy between the two sets of data is fairly uniform across a broad region of (low) temperatures and densities, including the standard condition (1 bar, 298.15K) that has been marked explicitly.
To get a feeling of the impact such an entropy error can have on the interpolation, Figure S8 shows the resulting EOS using the original (but incorrect) SCvH entropy data. Note that although we have selected an DFT energy offset to connect the free energy isotherms on both sides as smoothly as possible (see panel (c)), there still remains a noticeable mismatch. On the other hand, Fig. S8(d) shows the interpolated pressure is also apparently problematic. In particular, the wild wiggling behavior at the intermediate densities can even cause violation of the basic mechanical stability constraint [21] and hence definitely unphysical.
Notice that in Fig. S8(c), the free energy isotherms of SCvH appear to be wider than the ab initio region. Recalling the basic relation , the problematic interpolation can then be remedied by globally reducing the SCvH entropy to match the correct experimental value. This leads to the final EOS shown in Fig. 3 of the main text. Note we have also added an extra entropy offset of on the DFT side to make the interpolation look slightly better. Nevertheless, this residual error is much smaller than the SCvH entropy error discussed above and does not affect our essential arguments on the accuracy and reliability of the new EOS construction approach.
In principle, the remaining small error near the interpolation boundaries can be eliminated by using more accurate chemical models at low densities or electron structure methods (such as QMC) at high densities. Besides, the nuclear quantum effect may also have a slight impact on the ab initio entropy at low temperatures, which has been ignored in this work. This can be remedied either by adding some semi-empirical corrections [78, 79] or resorting to more accurate (and also more costly) path integral MD simulations [80].
Appendix G More details of comparison between different Jupiter’s adiabats
Figure S9 shows some more details of Jupiter’s adiabats calculated using various hydrogen EOSs. Compared to Fig. 4 of the main text, there is additionally the adiabat derived from the CMS19 EOS [11] (yellow line), which is clearly smoother in the ab initio region than both our result (red dashed line) and MH13 (blue line). This difference originates from the significant, non-monotonic slope change of our entropy isotherms, as shown in Fig. 1(a) of the main text. In Fig. S9, we also directly illustrate the entropy of our EOS by colored regions. The change of slope is again clearly evident, indicating that our EOS better captures the pressure-induced molecular dissociation than CMS19.
By definition, our calculated adiabat (red dashed line) follows the shape of a certain “stripe” of constant entropy, as clearly shown in Fig. S9. Note that the slightly imperfect behavior of our adiabat between 1 and 10GPa corresponds to the interpolation region of our EOS; see the previous section for more discussions. However, it is worth emphasizing that this remaining error will not affect the accuracy and reliability of our adiabat at higher pressures, where the disagreements between various adiabats make the biggest difference on resulting planetary model predictions [4, 7].
In Fig. S9, our adiabat is confined to the ab initio region explored in this work, specifically up to K and . It is in principle just a matter of additional DFT-MD simulations or interpolations with other theories [21, 81, 82, 83] to extend the hydrogen EOS to higher temperatures and densities. This will enable us to accurately study Jupiter’s deeper interior beyond the red dashed line in Fig. S9, or other astrophysical objects under even more extreme conditions, such as solar-type stars.

### v2

@@
 Introduction.
-The equation of state (EOS) of dense hydrogen (H) is one of the core ingredients to model the internal structure of giant gaseous planets, such as Jupiter and Saturn [1].
-With the wealth of data from recent spacecraft missions, Galileo [2] and Juno [3], which measured Jupiter’s atmospheric composition and gravitational field, lack of accurate knowledge of the thermodynamic properties of hydrogen is now considered as one of the major roadblocks in planetary science [4, 5, 6, 7].
-As parameterized planetary models already require several hypotheses [5, 8, 9, 10], an uncontested EOS is needed to advance the field.
-Since the high temperature and pressure inside these planets are still largely beyond the reach of current experimental techniques, one usually relies instead on numerical simulations, typically ab initio molecular dynamics (MD) based on electron structure methods such as density functional theory (DFT) or quantum Monte Carlo (QMC).
-Up to now, there have been several widely used ab initio EOSs for planetary science applications, including the one by Chabrier, Mazevet and Soubiran (CMS19) [11], by Militzer and Hubbard (MH13) [12], and the Rostock EOS (REOS2 [13], REOS3 [14]) of various versions.
-All of them are derived from DFT-MD simulations using the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) [15].
-Despite this common origin, however, they can make significant differences in the resulting Jupiter model, such as the location of adiabat and the predicted size of core mass [4, 7].
-The key factor behind these discrepancies of planetary models is the calculation of entropy, which has been recognized and discussed to some extent in several studies [12, 4, 6].
+The equation of state (EOS) of dense hydrogen (H) is one of the core ingredients to model the internal structure of giant gaseous planets, such as Jupiter and Saturn Helled et al.
+(2020).
+With the wealth of data from recent spacecraft missions, Galileo Johnson et al.
+(1992) and Juno Bolton et al.
+(2017), which measured Jupiter’s atmospheric composition and gravitational field, the lack of accurate knowledge of the thermodynamic properties of hydrogen is now considered as one of the major roadblocks in planetary science Miguel, Y.
+et al.
+(2016); Debras and Chabrier (2019); Mazevet, S.
+et al.
+(2022); Howard, S.
+et al.
+(2023).
+As parameterized planetary models already require several hypotheses Debras and Chabrier (2019); Nettelmann et al.
+(2021); Miguel, Y.
+et al.
+(2022); Militzer et al.
+(2022), an uncontested EOS is needed to advance the field.
+Since high-pressure experiments are difficult to perform and still cannot yield precise results, one usually relies instead on numerical simulations, typically ab initio molecular dynamics (MD) based on electron structure methods such as density functional theory (DFT) or quantum Monte Carlo (QMC).
+Up to now, there have been several widely used ab initio EOSs for planetary science applications, including the one by Chabrier, Mazevet and Soubiran (CMS19) Chabrier et al.
+(2019), by Militzer and Hubbard (MH13) Militzer and Hubbard (2013), and the Rostock EOS (REOS2 Nettelmann et al.
+(2012) and REOS3 Becker et al.
+(2014)) of various versions.
+All of them are derived from DFT-MD simulations using the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) Perdew et al.
+(1996).
+Despite this common origin, however, they can make significant differences in the resulting Jupiter model, such as the location of adiabat and the predicted size of core mass Miguel, Y.
+et al.
+(2016); Howard, S.
+et al.
+(2023).
+The key factor behind these discrepancies of planetary models is the calculation of entropy, which has been recognized and discussed to some extent in several studies Militzer and Hubbard (2013); Miguel, Y.
+et al.
+(2016); Mazevet, S.
+et al.
+(2022).
 Unlike observables such as energy and pressure , the Helmholtz free energy and entropy are not directly accessible from ab initio simulations and often rely on extra procedures called thermodynamic integration (TI).
-One common implementation of this technique involves interpolating the available data to create a continuous path on the temperature-density phase diagram [13, 4].
+One common implementation of this technique involves interpolating the available data to create a continuous path on the temperature-density phase diagram Nettelmann et al.
+(2012); Miguel, Y.
+et al.
+(2016).
 From basic statistical mechanics we have
 These two equations correspond to integrating the energy (pressure) of the system along isochores (isotherms), respectively.
-However, this approach has been criticized because, due to limited number of simulation points in practice, the interpolation procedure may introduce uncontrollable systematic errors [12].
-An alternative TI scheme is called the coupling constant integration (CCI) [16, 17] or Hamiltonian thermodynamic integration [18].
-By connecting the target state to an artificial Hamiltonian with tractable thermodynamic properties, this method further allows for the calculation of absolute entropy.
-However, it is also much more expensive due to the need to perform additional MD simulations, and therefore only suitable for investigating a small region of the phase diagram [19, 12].
-Besides, one has to carefully choose relevant parameters in the artificial system and MD simulations to ensure well-converged results [20], and the errors associated with discrete simulation points still remain.
+However, this approach has been criticized Militzer and Hubbard (2013) because, due to limited number of MD simulation points and inevitable statistical uncertainties in practice, the interpolation procedure may introduce systematic errors that are difficult to detect.
+An alternative TI scheme is called the coupling constant integration (CCI) Kirkwood (1935); de Wijs et al.
+(1998) or Hamiltonian thermodynamic integration Frenkel and Smit (2023).
+By connecting the target state to an artificial Hamiltonian with known thermodynamic properties, this method further allows for the calculation of absolute entropy.
+However, it is also much more expensive due to the need to perform additional MD simulations and therefore only suitable for investigating a small region of the phase diagram Morales et al.
+(2010a); Militzer and Hubbard (2013).
+Besides, one has to carefully choose relevant parameters in the artificial system and MD simulations to ensure well-converged results Bergermann et al.
+(2024), and the errors associated with discreteness and statistical fluctuation of the simulation points still remain.
 The uncertainties involved in entropy calculation are further exacerbated by the fact that the EOS used for planetary modeling is usually composed of several parts, each based on different theoretical approach and has different phase region of validity.
 For example, although DFT-MD can provide a good description of electron correlations at intermediate temperatures and densities, especially near pressure-induced dissociation, it will quickly become inefficient at lower densities .
-For such a region deep within the molecular phase, empirical chemical models such as the well known Saumon–Chabrier–van Horn (SCvH) EOS [21] are generally believed to be accurate and reliable.
-In practice, the data from these different methods needs to be connected in some way.
+For such a region deep within the molecular phase, empirical chemical models, such as the well-known Saumon–Chabrier–van Horn (SCvH) EOS Saumon et al.
+(1995) are generally believed to be accurate and reliable.
+In practice, the data from these different methods need to be connected in some way.
 This, however, could possibly introduce critical errors, as we will demonstrate below.
-In this Letter, we focus on the crucial issue of entropy calculation at the DFT-PBE level of theory, aiming to resolve long-standing disagreements among various Jupiter model predictions, which are highly sensitive to small changes in entropy [12, 4, 22, 6].
-We rely solely on ab initio theory and well-established experimental data without invoking indirect observations.
-The purpose of this paper is not to reconcile planetary models with spacecraft measurements [1, 5], but rather to provide a non-parametric hydrogen EOS, ranging from roughly 1 bar to 700GPa, that forms a solid foundation for future model revisions.
-We calculate the free energy and entropy of dense hydrogen using a recently developed method called flow matching [23, 24, 25].
-Unlike the demonstrative example shown in Ref.
-[26], here the calculations are based on DFT-MD simulations and thus have ab initio accuracy.
-Basically, we employ a machine-learned invertible transformation , which acts on configuration samples of the proton coordinates, to directly connect two phase points of interest.
-Given the transformation can bring the two states close enough to ensure a sufficient overlap, their free energy difference (scaled by temperature) can then be efficiently estimated as follows:
-where are the corresponding Boltzmann distributions, , being the Boltzmann constant, and
-are known as the forward and reverse work, respectively [27].
-In practice, the two estimators in Eq.
-(2) provide both upper and lower bound of the free energy difference, which can be pretty tight [27].
-In this way, one can accurately estimate the entropy difference based solely on samples and energy functions of the two end states, thus avoid errors arising from any kinds of interpolation between them as done in TI.
-Moreover, it is also fairly straightforward to compute the absolute entropy by just choosing one of the state to have tractable thermodynamic properties.
-The computational framework of Eqs.
-(2) and (3) is usually known as targeted free energy perturbation (TFEP) [28, 29].
-Thanks to the rapid development of deep generative models in recent years, the desired invertible map can now be efficiently represented by a class of generative models called normalizing flow [30], making TFEP emerge as a promising tool for free energy calculation [31, 32, 33, 34].
-In contrast to the “variational” approach, which has previously been applied to a broad range of many-body problems [35, 36, 32, 33, 37, 38, 39], the flow matching method we employ in this work represents a different training scheme of the flow model [23, 24, 25], which offers a variety of advantages in both accuracy and efficiency; see Supplemental Material [27] for details.
+In this Letter, we focus on the crucial issues of entropy calculation outlined above at the DFT-PBE level of theory, aiming to resolve long-standing disagreements among various Jupiter model predictions, which are highly sensitive to small changes in entropy Militzer and Hubbard (2013); Miguel, Y.
+et al.
+(2016, 2018); Mazevet, S.
+et al.
+(2022).
+The work is structured as follows: First, we calculate the free energy and entropy in the DFT-MD region using a recently developed method called flow matching Albergo and Vanden-Eijnden (2023); Lipman et al.
+(2023); Liu et al.
+(2023).
+As discussed below, such an independent approach is highly desired to benchmark the accuracy of TI and identify crucial deficiencies in previous works.
+Second, (and more importantly,) we devise a simple and efficient protocol to construct an EOS that can achieve excellent thermodynamic consistency across various theories and a broad range of conditions.
+Our procedure is primarily designed to provide a reliable and uncontested input for planetary modeling, thereby forming a solid foundation for future model refinements against spacecraft measurements and further applicability to other materials Helled et al.
+(2020); Debras and Chabrier (2019).
+Flow matching.
+Consider any two given phase points and of the hydrogen system.
+When calculating their free energy difference using TI, one needs to create a continuous path between them by interpolation.
+On the other hand, flow matching is built on the framework of targeted free energy perturbation (TFEP) Jarzynski (2002); Hahn and Then (2009), which solely relies on information about the two end states 0 and 1; see the Eqs.
+(2) and (3) below.
+As a result, it can avoid any errors arising from a manual interpolation procedure and thus serves as a valuable benchmark for TI.
+More specifically, within the framework of TFEP, the free energy difference (scaled by temperature) between the two states can be written as the following estimators:
+Here, denotes a configuration sample of the proton coordinates, and are the corresponding Boltzmann distributions, with , being the Boltzmann constant. The quantities
+are known as the forward and reverse work, respectively SM .
+Note that the above equations hold in principle for any invertible transformation that acts on the configuration space.
+In particular, when is the identity map, Eqs.
+(2) and (3) reduce to the usual formulas for importance sampling.
+Moreover, we note that by simply selecting one of the states to have known thermodynamic properties, it is also fairly straightforward to compute the absolute free energy and entropy of the system.
+In practice, due to the limited number of samples, Eq.
+(2) can be efficiently estimated with low variance only for some carefully designed transformation , which should bring the two states of interest close enough to ensure sufficient overlap of their probability densities.
+In this work, we achieve this goal by representing as a class of generative neural network called normalizing flow Papamakarios et al.
+(2021) and training the network using the flow matching technique Albergo and Vanden-Eijnden (2023); Lipman et al.
+(2023); Liu et al.
+(2023).
+Note that unlike the demonstrative example shown in Ref.
+Zhao and Wang (2023), here the calculations are based on DFT-MD simulations and thus have ab initio accuracy.
+A more detailed introduction to the relevant background is provided in the Supplemental Material SM .
 Entropy from ab initio data.
-We employ the flow matching method to study the ab initio portion of the phase diagram, with K and .
+We employ the flow matching method to study the ab initio portion of the phase diagram with K and .
 Specifically, we repeatedly train a transformation and use Eq.
-(2) to compute the entropy difference for each neighboring pair of states on a grid; see Supplemental Material [27] for an example.
+(2) to compute the entropy difference for each neighboring pair of states on a grid; see Supplemental Material SM for an example.
 We also calculate the absolute entropy of a reference state within the same framework and then shift the entropy of other states rigidly.
-In practice, we choose the reference point at and [27], where the computed value of the absolute entropy is in excellent agreement with those reported by Morales et al.
-[19] based on coupled electron-ion Monte Carlo (CEIMC) simulations.
-Figure 1 shows our results (solid line) for several isotherms of the specific entropy and pressure , together with two widely used hydrogen EOSs in past literature: CMS19 [11] (dotted line) and REOS3 [14] (dashed line).
-For convenience, we also include an error bar of absolute entropy at the reference point, and the CEIMC entropy of Morales et al.
-[19] (dash-dotted line) at relatively high densities.
-Note the REOS3 entropy is calculated by us using TI (see Eq.
-(1)), and, for sake of visualization, we have introduced a global entropy offset to make it match our data at the bottom-left phase point, i.e., and .
-We do not compare with the MH13 EOS [12] since it is derived directly for hydrogen-helium mixtures.
+In practice, we choose the reference point at and SM , where the computed value of the absolute entropy is in excellent agreement with those reported by Morales et al.
+Morales et al.
+(2010a) based on coupled electron-ion Monte Carlo (CEIMC) simulations.
+Figure 1 shows our results (solid line) for several isotherms of the specific entropy and pressure , together with two widely used hydrogen EOSs in past literature: CMS19 Chabrier et al.
+(2019) (dotted line) and REOS3 Becker et al.
+(2014) (dashed line).
+For convenience, we also include an error bar of absolute entropy at the reference point and the CEIMC entropy of Morales et al.
+Morales et al.
+(2010a) (dash-dotted line) at relatively high densities.
+Note that the REOS3 entropy is calculated by us using TI (see Eq.
+(1)), and, for the sake of visualization, we have introduced a global entropy offset to make it match our data at the bottom-left phase point, i.e., and .
+We do not compare with the MH13 EOS Militzer and Hubbard (2013) since it was derived directly for a hydrogen-helium mixture, rather than the pure hydrogen system studied here.
 One striking feature of Fig.
-1(a) is the non-monotonic behavior of entropy at low temperatures in both our results and REOS3, which is qualitatively different from CMS19.
+1(a) is the nonmonotonic behavior of entropy at low temperatures in both our results and REOS3, which is qualitatively different from CMS19.
 Moreover, CMS19 does not feature the typical slope change of the relation, as shown in Fig.
-1(b), which is emanating from the critical point of the liquid-liquid phase transition between molecular and atomic phase [40, 41].
-This indicates that although derived from DFT-PBE data in the phase region considered here [11, 42, 43, 19], the CMS19 EOS does not closely capture the pressure-induced molecular dissociation, which could have a notable impact on Jupiter models [27].
-The observed behavior of entropy and pressure in our results, as described above, can be understood by virtue of the thermodynamic relation [21]: .
-In fact, by examining the pressure isochores as illustrated in the inset of Fig.
-1(b), the region where the pressure decreases with temperature corresponds precisely to the region of increasing entropy in Fig.
-1(a).
-Physically, they are just different manifestations of the molecular dissociation effects: the attractive force from broken molecular bonds can dominates the repulsion from kinetic motion, while the larger number of configurations involving only individual H atoms (compared to those with bound H2-like pairs) can compensate for the smaller occupied volume caused by increasing density.
+1(b), which is emanating from the critical point of the liquid-liquid phase transition between molecular and atomic phase Lorenzen et al.
+(2010); Geng et al.
+(2019).
+This indicates that although derived from DFT-PBE data in the phase region considered here Chabrier et al.
+(2019); Caillabet et al.
+(2011); Holst et al.
+(2008); Morales et al.
+(2010a), the CMS19 EOS does not closely capture the pressure-induced molecular dissociation, which could have a notable impact on Jupiter models SM .
+According to the thermodynamic relation Saumon et al.
+(1995) , the region of “abnormally” increasing entropy in Fig.
+1(a) corresponds precisely to that of decreasing pressure with temperature, as illustrated by the isochores in the inset of Fig.
+1(b).
+Physically, they are just different manifestations of the molecular dissociation effects: The attractive force from broken molecular bonds can dominate the repulsion from kinetic motion, while the larger number of configurations involving only individual H atoms (compared to those with bound H2-like pairs) can compensate for the smaller occupied volume caused by increasing density.
 Benchmarking thermodynamic integration.
-We can use the flow matching entropy to benchmark the TI approach, which simply amounts to integrating Eq.
+We use the flow matching entropy in the ab initio region to benchmark the TI approach, which simply amounts to integrating Eq.
 (1) using our pressure and energy data.
@@
 Besides, they do not depend on the choice of integration path and hence show very good thermodynamic consistency.
-Contrary to the previous belief [12], our observations indicate that TI performs quite well for a discrete grid with spacings in the order of K and .
-When combined with the absolute entropy calculation on only a single referent point, as done above, we can then achieve the same level of accuracy as the “full” CCI approach in Ref.
-[12], but with far less computational costs.
-Note that compared to REOS3, we have added some additional simulation points at relatively low densities to improve the characterization of molecular dissociation.
-Nevertheless, TI is not fundamentally flawed even with the relatively sparse grid of REOS3 in the ab initio region, and the resulting entropy remains generally consistent with our data, as shown by the dashed lines in Fig.
+Our observations indicate that TI performs quite well for a discrete grid with spacings on the order of K and .
+When combined with the absolute entropy calculation on only a single reference point, as done above, we can then achieve the same level of accuracy as the “full” CCI approach in Ref.
+Militzer and Hubbard (2013) but with far less computational costs.
+Note that, compared to REOS3, we have selected a slightly denser grid of simulation points between and to improve the characterization of molecular dissociation.
+Nevertheless, the “discreteness” error of TI in the ab initio region of REOS3 is still under control, and the resulting entropy remains generally consistent with our data, as shown by the dashed lines in Fig.
 1(a).
 However, TI becomes truly problematic when we further extend the integration path to touch other regions based on inconsistent theories.
-We can quantitatively illustrate such inconsistency by performing the line integral along each closed local square loop of the tabular data, which should in principle vanish everywhere.
-However, for REOS3 we found large deviation from this ideal behavior between the ab initio region and the chemical model at lower densities [44, 45], as shown in Figure 2.
-This implies that the resulting entropy will be ambiguously defined and strongly depend on the choice of integration path and initial reference point; see Supplemental Material [27] for more details.
-One can then identify an entropy error of approximately [27], which yields qualitatively large differences on planetary models derived in the past using this method [13, 4].
-A thermodynamically consistent construction.
-We have shown that performing a single TI over an EOS containing various inconsistent theories can lead to significant and uncontrolled errors of the entropy.
-A better approach is to perform TI on each component (i.e., region where the same theory applies) of the EOS separately, and then interpolate the resulting free energy function [27].
+We can quantitatively illustrate such inconsistency by performing the line integral along each closed local square loop of the tabular data, which should, in principle, vanish everywhere.
+However, for REOS3, we found large deviation from this ideal behavior between the ab initio region and the chemical model at lower densities Juranek et al.
+(2002); Holst et al.
+(2007), as shown in Fig.
+2.
+This implies that the resulting entropy will be ambiguously defined and strongly depend on the choice of integration path and initial reference point; see Supplemental Material SM for more details.
+One can then identify an entropy error of approximately SM , which yields qualitatively large differences on planetary models derived in the past using this method Nettelmann et al.
+(2012); Miguel, Y.
+et al.
+(2016).
+Thermodynamically consistent construction.
+An inherent difficulty of the “single TI over the entire phase diagram” approach stems from the fact that energy and pressure are different partial derivatives of the same free energy function and hence not independent.
+Consequently, it turns out to be highly challenging to simultaneously interpolate both energy and pressure data produced by different theories in a consistent way, as demonstrated in Fig.
+2 for the case of REOS3.
+A better approach is to perform TI on each component (i.e., region where the same theory applies) of the EOS separately and then interpolate the resulting free energy function SM .
 In this way, one can derive the pressure and energy near the interpolation boundaries by taking partial derivatives, which are naturally consistent by construction.
@@
 This ensures that any residual interpolation errors that one can make at the boundaries remain localized and do not propagate elsewhere.
-Figure 3(a-d) show our final EOS (solid lines), which consist of the ab initio DFT-MD part at , the SCvH EOS at , and an interpolation region between them.
+Figures 3(a-d) show our final EOS (solid lines), which consists of the ab initio DFT-MD part at (only above K), the SCvH EOS at (covering a broader temperature range down to K), and an interpolation region between them.
 For comparison, we also include the entropy and pressure isotherms of Miguel et al.
-[4] (dashed lines), which are obtained by a single TI over the entire phase region of REOS3.
-Note the new interpolation procedure depends very sensitively on the entropy difference between SCvH and the ab initio data [27].
-In doing this, we then surprisingly found the reported (absolute) entropies of SCvH at room conditions (e.g., at standard pressure 1 bar and temperature 298.15K) are about (or ) higher than both the experimental data [46, 47] and theoretical calculations [48].
-Only after such errors are fixed (by applying a global entropy shift on the SCvH side) can we achieve a high-quality interpolation that respects fundamental thermodynamic constraints; see Supplemental Material [27] for details.
-This observation, which results from totally independent experimental or numerical benchmarks at both low ( 1 bar) and high ( 700GPa) pressures, is a very strong evidence on the accuracy of our construction.
-Fig.
-3(e) also shows the local loop integral values of our final EOS; compared to Fig.
-2, one clearly sees the new interpolation approach can indeed yield much better thermodynamic consistency over the entire phase diagram, as discussed above.
+Miguel, Y.
+et al.
+(2016) (dashed lines), which are obtained by a single TI over the entire phase region of REOS3.
+Note that the new interpolation procedure depends very sensitively on the entropy difference between SCvH and the ab initio data SM .
+In particular, a high-quality interpolation can only be achieved after globally subtracting a constant of from the originally reported SCvH entropy.
+Remarkably, this observation aligns precisely with various independent numerical Chabrier et al.
+(2019) and experimental references Cox et al.
+(1989); Lemmon et al.
+(2023).
+There, the entropy constant was recognized as , which accounts for the contribution of proton spin degrees of freedom that are implicitly neglected on the DFT-MD side Saumon and Chabrier ; see Supplemental Material SM for details.
+This strongly indicates that our new protocol is indeed the very approach to construct accurate and uncontested hydrogen EOS for planetary modeling.
+Figure 3(e) also shows the local loop integral values of our final EOS; compared to Fig.
+2, one clearly sees that the new protocol can indeed yield much better thermodynamic consistency over the entire phase diagram, as discussed above.
 Jupiter’s adiabat.
 We demonstrate the utility of the newly constructed hydrogen EOS by performing a preliminary calculation of Jupiter’s adiabat.
-We assume a homogeneous model with no compact core, and set a helium mass fraction of to ensure a fair comparison with previous works [4, 22, 12].
+We assume a homogeneous model with no compact core and set a helium mass fraction of to ensure a fair comparison with previous works Miguel, Y.
+et al.
+(2016, 2018); Militzer and Hubbard (2013).
 We adopt a simple linear mixing approximation for the entropy of mixtures, using the REOS3 EOS of helium derived by Miguel et al.
 in Ref.
-[22].
-The starting point is bar and K [1], which corresponds to an entropy (for hydrogen) of .
-Figure 4 shows our result for the adiabat (red dashed line), together with those derived from the REOS3 (black line) [22] and MH13 EOS (blue line) [12], respectively.
-For a comparison also with CMS19, see the Supplemental Material [27].
-Note the hydrogen EOS used in the REOS3 adiabat was produced by Miguel et al.
-[4, 22] using a single TI over the entire phase region, as already shown by the dashed lines in Fig.
+Miguel, Y.
+et al.
+(2018).
+Figure 4 shows our result for the adiabat (red dashed line) together with those derived from the REOS3 (black line) Miguel, Y.
+et al.
+(2018) and MH13 EOS (blue line) Militzer and Hubbard (2013), respectively.
+For a comparison also with CMS19, see the Supplemental Material SM .
+Note that the hydrogen EOS used in the REOS3 adiabat was produced by Miguel et al.
+Miguel, Y.
+et al.
+(2016, 2018) using a single TI over the entire phase region, as already shown by the dashed lines in Fig.
 3.
@@
 3(a), which is in turn caused by the thermodynamic inconsistency shown in Fig.
-2 [27].
-Notice such a cooler adiabat passes very close to the miscibility boundary for hydrogen-helium mixtures, shown as the yellow line in Fig.
-4, albeit derived from DFT-MD simulations using an exchange-correlation functional different from PBE [49].
-This suggests an increased likelihood of phase separation between hydrogen and helium in Jupiter’s deep interior, causing the formation of so-called helium rain [1, 50, 51].
+2 SM .
+Notice that such a cooler adiabat passes very close to the miscibility boundary for hydrogen-helium mixtures, shown by the yellow line in Fig.
+4, albeit derived from DFT-MD simulations using an exchange-correlation functional different from PBE Schöttler and Redmer (2018).
+This suggests an increased likelihood of phase separation between hydrogen and helium in Jupiter’s deep interior, causing the formation of so-called helium rain Helled et al.
+(2020); Vorberger et al.
+(2007); Chang et al.
+(2024).
 On the other hand, our adiabat at high pressures is apparently closer to the MH13 prediction, as shown by the blue line in Fig.
@@
 However, according to our earlier benchmark using flow matching, this definitely does not imply that TI is inherently inferior to the CCI approach for a finite number of simulation points, as speculated in Ref.
-[12].
+Militzer and Hubbard (2013).
 Another observation from Fig.
-4 is that the MH13 adiabat [12] was derived only for the ab initio region at high pressures.
+4 is that the MH13 adiabat Militzer and Hubbard (2013) was derived only for the ab initio region at high pressures.
 In fact, Ref.
-[12] did not really attempt a thermodynamic consistent interpolation with other regions to construct a complete EOS, which, however, is definitely needed for a thorough study of planetary models.
+Militzer and Hubbard (2013) did not really attempt a thermodynamic consistent interpolation with other regions to construct a complete EOS, which, however, is definitely needed for a thorough study of planetary models.
 Summary and outlook.
-In this Letter, we closely examine the entropy calculation of dense hydrogen by benchmarking TI with an independent flow matching method.
-Based on this, we then construct a new hydrogen EOS that is both accurate and thermodynamically consistent over several orders of magnitude of pressure, which is crucial for reliable and conclusive planetary model predictions.
-Notably, the interpolation quality at the boundaries serves as a highly sensitive and stringent test for the accuracy of the theories on both sides [27].
-Indeed, the new approach has allowed us to identify and correct a significant discrepancy in SCvH relative to experimental data at low pressures, which is fairly surprising and remarkable given the chemical model’s long history of nearly 30 years.
-We conclusively address the long-standing issue of entropy disagreements among various popular EOSs used in planetary modeling [11, 12, 14], all of which are derived from the same DFT-PBE level of electronic structure theory.
-To further improve the hydrogen EOS and potentially reconcile models with observational data [5], one can upgrade the description of electronic correlation from DFT-PBE to a higher-level theory [52, 53, 54, 39].
-We also note that as a generic tool, flow matching can also be used to study other rich phases of hydrogen [55, 56] or other interesting condensed matter systems [51, 57], where accurate calculation of free energy plays a significant role.
+In this Letter, we closely examine the entropy calculation of dense hydrogen in the ab initio region by benchmarking TI with an independent flow matching method.
+Building on this, we then construct a new hydrogen EOS that is both accurate and thermodynamically consistent over several orders of magnitude of pressure (roughly 1bar 700GPa).
+Remarkably, the interpolation quality at the boundaries serves as a highly sensitive and stringent “detector” for the accuracy of the theories on both sides SM .
+Such a “self-correcting” mechanism of the new protocol is exactly what we need to ensure steady convergence toward a truly reliable and conclusive hydrogen EOS for planetary modeling.
+We conclusively address the long-standing issue of entropy disagreements among various popular EOSs used in planetary modeling Chabrier et al.
+(2019); Militzer and Hubbard (2013); Becker et al.
+(2014), all of which are derived from the same DFT-PBE level of electronic structure theory.
+To further improve the hydrogen EOS and potentially reconcile models with observational data Debras and Chabrier (2019), one can upgrade the description of electronic correlation from DFT-PBE to a higher-level theory Pierleoni et al.
+(2016); Mazzola et al.
+(2018); Nakano et al.
+(2020); Xie et al.
+(2023a).
+We also note that as a generic tool, flow matching can also be used to study other rich phases of hydrogen Morales et al.
+(2010b); Niu et al.
+(2023) or other interesting condensed matter systems Chang et al.
+(2024); Wang et al.
+(2023), where accurate calculation of free energy plays a significant role.
 Acknowledgements.
-We acknowledge useful discussions with Lei Wang, Ravit Helled, Gilles Chabrier, and Armin Bergermann.
+We acknowledge useful discussions with Lei Wang, Ravit Helled, Didier Saumon, Gilles Chabrier, and Armin Bergermann.
 We also acknowledge Cesare Cozza for providing initial data and support with the DFT calculations.
@@
 and G.M.
-acknowledge financial support from the Swiss National Science Foundation (grant PCEFP2_203455).
+acknowledge financial support from the Swiss National Science Foundation (Grant No.
+PCEFP2_203455).
 S.H.
-acknowledges financial support from the Swiss National Science Foundation (grant 200020_215634).References
-- Helled et al. [2020] R. Helled, G. Mazzola, and R. Redmer, Understanding dense hydrogen at planetary conditions, Nature Reviews Physics 2, 562 (2020).
-- Johnson et al. [1992] T. V. Johnson, C. M. Yeates, and R. Young, Space science reviews volume on Galileo Mission overview, Space Science Reviews 60, 3 (1992).
+also acknowledges financial support from the Swiss National Science Foundation (Grant No.
+200020_215634).References
+- Helled et al. (2020) Ravit Helled, Guglielmo Mazzola, and Ronald Redmer, “Understanding dense hydrogen at planetary conditions,” Nature Reviews Physics 2, 562–574 (2020).
+- Johnson et al. (1992) T. V. Johnson, C. M. Yeates, and R. Young, “Space science reviews volume on Galileo Mission overview,” Space Science Reviews 60, 3–21 (1992).
 - Bolton et al.
-[2017] S.
+(2017) S.
 J.
@@
 Thorne, and R.
-Thorpe, The Juno Mission, Space Science Reviews 213, 5 (2017).
-- Miguel, Y. et al. [2016] Miguel, Y., Guillot, T., and Fayon, L., Jupiter internal structure: the effect of different equations of state, A&A 596, A114 (2016).
-- Debras and Chabrier [2019] F. Debras and G. Chabrier, New Models of Jupiter in the Context of Juno and Galileo, The Astrophysical Journal 872, 100 (2019).
-- Mazevet, S. et al. [2022] Mazevet, S., Licari, A., and Soubiran, F., Benchmarking the ab initio hydrogen equation of state for the interior structure of jupiter, A&A 664, A112 (2022).
+Thorpe, “The Juno Mission,” Space Science Reviews 213, 5–37 (2017).
+- Miguel, Y. et al. (2016) Miguel, Y., Guillot, T., and Fayon, L., “Jupiter internal structure: the effect of different equations of state,” A&A 596, A114 (2016).
+- Debras and Chabrier (2019) Florian Debras and Gilles Chabrier, “New Models of Jupiter in the Context of Juno and Galileo,” The Astrophysical Journal 872, 100 (2019).
+- Mazevet, S. et al. (2022) Mazevet, S., Licari, A., and Soubiran, F., “Benchmarking the ab initio hydrogen equation of state for the interior structure of jupiter,” A&A 664, A112 (2022).
 - Howard, S.
 et al.
-[2023] Howard, S., Guillot, T., Bazot, M., Miguel, Y., Stevenson, D.
+(2023) Howard, S., Guillot, T., Bazot, M., Miguel, Y., Stevenson, D.
 J., Galanti, E., Kaspi, Y., Hubbard, W.
-B., Militzer, B., Helled, R., Nettelmann, N., Idini, B., and Bolton, S., Jupiter’s interior from Juno: Equation-of-state uncertainties and dilute core extent, A&A 672, A33 (2023).
+B., Militzer, B., Helled, R., Nettelmann, N., Idini, B., and Bolton, S., “Jupiter’s interior from Juno: Equation-of-state uncertainties and dilute core extent,” A&A 672, A33 (2023).
 - Nettelmann et al.
-[2021] N.
+(2021) N.
 Nettelmann, N.
@@
 Mankovich, and S.
-Bolton, Theory of Figures to the Seventh Order and the Interiors of Jupiter and Saturn, The Planetary Science Journal 2, 241 (2021).
+Bolton, “Theory of Figures to the Seventh Order and the Interiors of Jupiter and Saturn,” The Planetary Science Journal 2, 241 (2021).
 - Miguel, Y.
 et al.
-[2022] Miguel, Y., Bazot, M., Guillot, T., Howard, S., Galanti, E., Kaspi, Y., Hubbard, W.
+(2022) Miguel, Y., Bazot, M., Guillot, T., Howard, S., Galanti, E., Kaspi, Y., Hubbard, W.
 B., Militzer, B., Helled, R., Atreya, S.
@@
 P., Durante, D., Kulowski, L., Lunine, J.
-I., Stevenson, D., and Bolton, S., Jupiter’s inhomogeneous envelope, A&A 662, A18 (2022).
+I., Stevenson, D., and Bolton, S., “Jupiter’s inhomogeneous envelope,” A&A 662, A18 (2022).
 - Militzer et al.
-[2022] B.
-Militzer, W.
-B.
-Hubbard, S.
-Wahl, J.
-I.
-Lunine, E.
-Galanti, Y.
-Kaspi, Y.
-Miguel, T.
-Guillot, K.
-M.
-Moore, M.
-Parisi, J.
-E.
+(2022) Burkhard Militzer, William B.
+Hubbard, Sean Wahl, Jonathan I.
+Lunine, Eli Galanti, Yohai Kaspi, Yamila Miguel, Tristan Guillot, Kimberly M.
+Moore, Marzia Parisi, John E.
 P.
-Connerney, R.
-Helled, H.
-Cao, C.
-Mankovich, D.
-J.
-Stevenson, R.
-S.
-Park, M.
-Wong, S.
-K.
-Atreya, J.
-Anderson, and S.
-J.
-Bolton, Juno Spacecraft Measurements of Jupiter’s Gravity Imply a Dilute Core, The Planetary Science Journal 3, 185 (2022).
-- Chabrier et al. [2019] G. Chabrier, S. Mazevet, and F. Soubiran, A new equation of state for dense hydrogen–helium mixtures, The Astrophysical Journal 872, 51 (2019).
-- Militzer and Hubbard [2013] B.
+Connerney, Ravid Helled, Hao Cao, Christopher Mankovich, David J.
+Stevenson, Ryan S.
+Park, Mike Wong, Sushil K.
+Atreya, John Anderson, and Scott J.
+Bolton, “Juno Spacecraft Measurements of Jupiter’s Gravity Imply a Dilute Core,” The Planetary Science Journal 3, 185 (2022).
+- Chabrier et al. (2019) G. Chabrier, S. Mazevet, and F. Soubiran, “A new equation of state for dense hydrogen–helium mixtures,” The Astrophysical Journal 872, 51 (2019).
+- Militzer and Hubbard (2013) B.
 Militzer and W.
 B.
-Hubbard, Ab initio equation of state for hydrogen–helium mixtures with recalibration of the giant-planet mass–radius relation, The Astrophysical Journal 774, 148 (2013).
-- Nettelmann et al. [2012] N. Nettelmann, A. Becker, B. Holst, and R. Redmer, Jupiter models with improved ab initio hydrogen equation of state (h-reos.2), The Astrophysical Journal 750, 52 (2012).
+Hubbard, “Ab initio equation of state for hydrogen–helium mixtures with recalibration of the giant-planet mass–radius relation,” The Astrophysical Journal 774, 148 (2013).
+- Nettelmann et al. (2012) N. Nettelmann, A. Becker, B. Holst, and R. Redmer, “Jupiter models with improved ab initio hydrogen equation of state (h-reos.2),” The Astrophysical Journal 750, 52 (2012).
 - Becker et al.
-[2014] A.
-Becker, W.
-Lorenzen, J.
-J.
-Fortney, N.
-Nettelmann, M.
-Schöttler, and R.
-Redmer, Ab initio equations of state for hydrogen (h-reos.3) and helium (he-reos.3) and their implications for the interior of brown dwarfs, The Astrophysical Journal Supplement Series 215, 21 (2014).
-- Perdew et al. [1996] J. P. Perdew, K. Burke, and M. Ernzerhof, Generalized gradient approximation made simple, Phys. Rev. Lett. 77, 3865 (1996).
-- Kirkwood [1935] J. G. Kirkwood, Statistical Mechanics of Fluid Mixtures, The Journal of Chemical Physics 3, 300 (1935).
-- de Wijs et al. [1998] G. A. de Wijs, G. Kresse, and M. J. Gillan, First-order phase transitions by first-principles free-energy calculations: The melting of al, Phys. Rev. B 57, 8223 (1998).
-- Frenkel and Smit [2023] D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications, 3rd ed. (Academic Press, 2023).
-- Morales et al. [2010a] M. A. Morales, C. Pierleoni, and D. M. Ceperley, Equation of state of metallic hydrogen from coupled electron-ion monte carlo simulations, Phys. Rev. E 81, 021202 (2010a).
-- Bergermann et al. [2024] A. Bergermann, M. French, and R. Redmer, Ab initio calculation of the miscibility diagram for mixtures of hydrogen and water, Phys. Rev. B 109, 174107 (2024).
-- Saumon et al. [1995] D. Saumon, G. Chabrier, and H. M. van Horn, An Equation of State for Low-Mass Stars and Giant Planets, Astrophysical Journal Supplement 99, 713 (1995).
-- Miguel, Y. et al. [2018] Miguel, Y., Guillot, T., and Fayon, L., Jupiter internal structure: the effect of different equations of state (corrigendum), A&A 618, C2 (2018).
-- Albergo and Vanden-Eijnden [2023] M.
-S.
-Albergo and E.
-Vanden-Eijnden, Building normalizing flows with stochastic interpolants, in The Eleventh International Conference on Learning Representations (2023).
-- Lipman et al. [2023] Y. Lipman, R. T. Q. Chen, H. Ben-Hamu, M. Nickel, and M. Le, Flow matching for generative modeling, in The Eleventh International Conference on Learning Representations (2023).
+(2014) Andreas Becker, Winfried Lorenzen, Jonathan J.
+Fortney, Nadine Nettelmann, Manuel Schöttler, and Ronald Redmer, “Ab initio equations of state for hydrogen (h-reos.3) and helium (he-reos.3) and their implications for the interior of brown dwarfs,” The Astrophysical Journal Supplement Series 215, 21 (2014).
+- Perdew et al. (1996) John P. Perdew, Kieron Burke, and Matthias Ernzerhof, “Generalized gradient approximation made simple,” Phys. Rev. Lett. 77, 3865–3868 (1996).
+- Kirkwood (1935) John G. Kirkwood, “Statistical Mechanics of Fluid Mixtures,” The Journal of Chemical Physics 3, 300–313 (1935).
+- de Wijs et al.
+(1998) Gilles A.
+de Wijs, Georg Kresse, and Michael J.
+Gillan, “First-order phase transitions by first-principles free-energy calculations: The melting of al,” Phys.
+Rev.
+B 57, 8223–8234 (1998).
+- Frenkel and Smit (2023) D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications, 3rd ed. (Academic Press, 2023).
+- Morales et al.
+(2010a) Miguel A.
+Morales, Carlo Pierleoni, and D.
+M.
+Ceperley, “Equation of state of metallic hydrogen from coupled electron-ion monte carlo simulations,” Phys.
+Rev.
+E 81, 021202 (2010a).
+- Bergermann et al. (2024) Armin Bergermann, Martin French, and Ronald Redmer, “Ab initio calculation of the miscibility diagram for mixtures of hydrogen and water,” Phys. Rev. B 109, 174107 (2024).
+- Saumon et al. (1995) D. Saumon, G. Chabrier, and H. M. van Horn, “An Equation of State for Low-Mass Stars and Giant Planets,” Astrophysical Journal Supplement 99, 713 (1995).
+- Miguel, Y. et al. (2018) Miguel, Y., Guillot, T., and Fayon, L., “Jupiter internal structure: the effect of different equations of state (corrigendum),” A&A 618, C2 (2018).
+- Albergo and Vanden-Eijnden (2023) Michael Samuel Albergo and Eric Vanden-Eijnden, “Building normalizing flows with stochastic interpolants,” in The Eleventh International Conference on Learning Representations (2023).
+- Lipman et al.
+(2023) Yaron Lipman, Ricky T.
+Q.
+Chen, Heli Ben-Hamu, Maximilian Nickel, and Matthew Le, “Flow matching for generative modeling,” in The Eleventh International Conference on Learning Representations (2023).
 - Liu et al.
-[2023] X.
-Liu, C.
-Gong, and qiang liu, Flow straight and fast: Learning to generate and transfer data with rectified flow, in The Eleventh International Conference on Learning Representations (2023).
-- Zhao and Wang [2023] L. Zhao and L. Wang, Bounding free energy difference with flow matching, Chinese Physics Letters 40, 120201 (2023).
-- [27] See Supplemental Material for more details on the theoretical formulations of targeted free energy perturbation and flow matching, as well as their implementation and some illustrative examples in practice.
+(2023) Xingchao Liu, Chengyue Gong, and qiang liu, “Flow straight and fast: Learning to generate and transfer data with rectified flow,” in The Eleventh International Conference on Learning Representations (2023).
+- Jarzynski (2002) C. Jarzynski, “Targeted free energy perturbation,” Phys. Rev. E 65, 046122 (2002).
+- Hahn and Then (2009) A. M. Hahn and H. Then, “Using bijective maps to improve free-energy estimates,” Phys. Rev. E 79, 011113 (2009).
+- (28) See Supplemental Material for more details on the theoretical formulations of targeted free energy perturbation and flow matching, as well as their implementation and some illustrative examples in practice.
 There are also further discussions about the effects of thermodynamic inconsistency on entropy calculation, our new EOS construction approach, and the comparison between Jupiter’s adiabats derived from various hydrogen EOSs.
-- Jarzynski [2002] C. Jarzynski, Targeted free energy perturbation, Phys. Rev. E 65, 046122 (2002).
-- Hahn and Then [2009] A. M. Hahn and H. Then, Using bijective maps to improve free-energy estimates, Phys. Rev. E 79, 011113 (2009).
 - Papamakarios et al.
-[2021] G.
-Papamakarios, E.
-Nalisnick, D.
-J.
-Rezende, S.
-Mohamed, and B.
-Lakshminarayanan, Normalizing flows for probabilistic modeling and inference, Journal of Machine Learning Research 22, 1 (2021).
-- Wirnsberger et al.
-[2020] P.
-Wirnsberger, A.
-J.
-Ballard, G.
-Papamakarios, S.
-Abercrombie, S.
-Racanière, A.
-Pritzel, D.
-Jimenez Rezende, and C.
-Blundell, Targeted free energy estimation via learned mappings, The Journal of Chemical Physics 153, 144112 (2020).
-- Wirnsberger et al.
-[2022] P.
-Wirnsberger, G.
-Papamakarios, B.
-Ibarz, S.
-Racanière, A.
-J.
-Ballard, A.
-Pritzel, and C.
-Blundell, Normalizing flows for atomic solids, Machine Learning: Science and Technology 3, 025009 (2022).
-- Ahmad and Cai [2022] R. Ahmad and W. Cai, Free energy calculation of crystalline solids using normalizing flows, Modelling and Simulation in Materials Science and Engineering 30, 065007 (2022).
-- Molina-Taborda et al.
-[2024] A.
-Molina-Taborda, P.
-Cossio, O.
-Lopez-Acevedo, and M.
-Gabrié, Active Learning of Boltzmann Samplers and Potential Energies with Quantum Mechanical Accuracy, Journal of Chemical Theory and Computation 20, 8833 (2024).
-- Li and Wang [2018] S.-H. Li and L. Wang, Neural network renormalization group, Phys. Rev. Lett. 121, 260601 (2018).
-- Wu et al. [2019] D. Wu, L. Wang, and P. Zhang, Solving statistical mechanics using variational autoregressive networks, Phys. Rev. Lett. 122, 080602 (2019).
-- Xie et al. [2022] H. Xie, L. Zhang, and L. Wang, Ab-initio study of interacting fermions at finite temperature with neural canonical transformation, Journal of Machine Learning 1, 38 (2022).
-- Xie et al. [2023a] H. Xie, L. Zhang, and L. Wang, of two-dimensional electron gas: A neural canonical transformation study, SciPost Phys. 14, 154 (2023a).
-- Xie et al. [2023b] H. Xie, Z.-H. Li, H. Wang, L. Zhang, and L. Wang, Deep variational free energy approach to dense hydrogen, Phys. Rev. Lett. 131, 126501 (2023b).
-- Lorenzen et al. [2010] W. Lorenzen, B. Holst, and R. Redmer, First-order liquid-liquid phase transition in dense hydrogen, Phys. Rev. B 82, 195107 (2010).
+(2021) George Papamakarios, Eric Nalisnick, Danilo Jimenez Rezende, Shakir Mohamed, and Balaji Lakshminarayanan, “Normalizing flows for probabilistic modeling and inference,” Journal of Machine Learning Research 22, 1–64 (2021).
+- Zhao and Wang (2023) Lu Zhao and Lei Wang, “Bounding free energy difference with flow matching,” Chinese Physics Letters 40, 120201 (2023).
+- Lorenzen et al. (2010) Winfried Lorenzen, Bastian Holst, and Ronald Redmer, “First-order liquid-liquid phase transition in dense hydrogen,” Phys. Rev. B 82, 195107 (2010).
 - Geng et al.
-[2019] H.
-Y.
+(2019) Hua Y.
 Geng, Q.
-Wu, M.
-Marqués, and G.
-J.
-Ackland, Thermodynamic anomalies and three distinct liquid-liquid transitions in warm dense liquid hydrogen, Phys.
+Wu, Miriam Marqués, and Graeme J.
+Ackland, “Thermodynamic anomalies and three distinct liquid-liquid transitions in warm dense liquid hydrogen,” Phys.
 Rev.
@@
 - Caillabet et al.
-[2011] L.
+(2011) L.
 Caillabet, S.
 Mazevet, and P.
-Loubeyre, Multiphase equation of state of hydrogen from ab initio calculations in the range 0.2 to 5 g/cc up to 10 ev, Phys.
+Loubeyre, “Multiphase equation of state of hydrogen from ab initio calculations in the range 0.2 to 5 g/cc up to 10 ev,” Phys.
 Rev.
 B 83, 094101 (2011).
-- Holst et al. [2008] B. Holst, R. Redmer, and M. P. Desjarlais, Thermophysical properties of warm dense hydrogen using quantum molecular dynamics simulations, Phys. Rev. B 77, 184201 (2008).
+- Holst et al.
+(2008) Bastian Holst, Ronald Redmer, and Michael P.
+Desjarlais, “Thermophysical properties of warm dense hydrogen using quantum molecular dynamics simulations,” Phys.
+Rev.
+B 77, 184201 (2008).
 - Juranek et al.
-[2002] H.
-Juranek, R.
-Redmer, and Y.
-Rosenfeld, Fluid variational theory for pressure dissociation in dense hydrogen: Multicomponent reference system and nonadditivity effects, The Journal of Chemical Physics 117, 1768 (2002).
-- Holst et al. [2007] B. Holst, N. Nettelmann, and R. Redmer, Equation of state for dense hydrogen and plasma phase transition, Contributions to Plasma Physics 47, 368 (2007).
-- Cox et al. [1989] J. D. Cox, D. D. Wagman, and V. A. Medvedev, Codata key values for thermodynamics, in Series on Thermodynamic Properties (1989).
+(2002) Hauke Juranek, Ronald Redmer, and Yaakov Rosenfeld, “Fluid variational theory for pressure dissociation in dense hydrogen: Multicomponent reference system and nonadditivity effects,” The Journal of Chemical Physics 117, 1768–1774 (2002).
+- Holst et al. (2007) B. Holst, N. Nettelmann, and R. Redmer, “Equation of state for dense hydrogen and plasma phase transition,” Contributions to Plasma Physics 47, 368–374 (2007).
+- Cox et al. (1989) John D Cox, Donald D Wagman, and Vadim Andreevich Medvedev, “Codata key values for thermodynamics,” in Series on Thermodynamic Properties (1989).
 - Lemmon et al.
-[2023] E.
-W.
-Lemmon, I.
-H.
-Bell, M.
-L.
-Huber, and M.
-O.
-McLinden, Thermophysical properties of fluid systems, in NIST Chemistry WebBook, NIST Standard Reference Database Number 69, edited by P.
-Linstrom and W.
+(2023) Eric W.
+Lemmon, Ian H.
+Bell, Marcia L.
+Huber, and Mark O.
+McLinden, “Thermophysical properties of fluid systems,” in NIST Chemistry WebBook, NIST Standard Reference Database Number 69, edited by P.J.
+Linstrom and W.G.
 Mallard (National Institute of Standards and Technology, Gaithersburg MD, 20899, 2023).
-- NIS [2022] NIST Computational Chemistry Comparison and Benchmark Database, NIST Standard Reference Database Number 101 (2022).
-- Schöttler and Redmer [2018] M. Schöttler and R. Redmer, Ab initio calculation of the miscibility diagram for hydrogen-helium mixtures, Phys. Rev. Lett. 120, 115703 (2018).
-- Vorberger et al. [2007] J. Vorberger, I. Tamblyn, B. Militzer, and S. A. Bonev, Hydrogen-helium mixtures in the interiors of giant planets, Phys. Rev. B 75, 024206 (2007).
+- (39) Didier Saumon and Gilles Chabrier, private communication.
+- Schöttler and Redmer (2018) Manuel Schöttler and Ronald Redmer, “Ab initio calculation of the miscibility diagram for hydrogen-helium mixtures,” Phys. Rev. Lett. 120, 115703 (2018).
+- Vorberger et al. (2007) J. Vorberger, I. Tamblyn, B. Militzer, and S. A. Bonev, “Hydrogen-helium mixtures in the interiors of giant planets,” Phys. Rev. B 75, 024206 (2007).
 - Chang et al.
-[2024] X.
-Chang, B.
-Chen, Q.
-Zeng, H.
-Wang, K.
-Chen, Q.
-Tong, X.
-Yu, D.
-Kang, S.
-Zhang, F.
-Guo, Y.
-Hou, Z.
-Zhao, Y.
-Yao, Y.
-Ma, and J.
-Dai, Theoretical evidence of h-he demixing under jupiter and saturn conditions, Nature Communications 15, 8543 (2024).
+(2024) Xiaoju Chang, Bo Chen, Qiyu Zeng, Han Wang, Kaiguo Chen, Qunchao Tong, Xiaoxiang Yu, Dongdong Kang, Shen Zhang, Fangyu Guo, Yong Hou, Zengxiu Zhao, Yansun Yao, Yanming Ma, and Jiayu Dai, “Theoretical evidence of h-he demixing under jupiter and saturn conditions,” Nature Communications 15, 8543 (2024).
 - Pierleoni et al.
-[2016] C.
-Pierleoni, M.
-A.
-Morales, G.
-Rillo, M.
-Holzmann, and D.
-M.
-Ceperley, Liquid–liquid phase transition in hydrogen by coupled electron–ion monte carlo simulations, Proceedings of the National Academy of Sciences 113, 4953 (2016).
+(2016) Carlo Pierleoni, Miguel A.
+Morales, Giovanni Rillo, Markus Holzmann, and David M.
+Ceperley, “Liquid–liquid phase transition in hydrogen by coupled electron–ion monte carlo simulations,” Proceedings of the National Academy of Sciences 113, 4953–4957 (2016).
 - Mazzola et al.
-[2018] G.
-Mazzola, R.
-Helled, and S.
-Sorella, Phase diagram of hydrogen and a hydrogen-helium mixture at planetary conditions by quantum monte carlo simulations, Phys.
+(2018) Guglielmo Mazzola, Ravit Helled, and Sandro Sorella, “Phase diagram of hydrogen and a hydrogen-helium mixture at planetary conditions by quantum monte carlo simulations,” Phys.
 Rev.
@@
 - Nakano et al.
-[2020] K.
-Nakano, C.
-Attaccalite, M.
-Barborini, L.
-Capriotti, M.
-Casula, E.
-Coccia, M.
-Dagrada, C.
-Genovese, Y.
-Luo, G.
-Mazzola, A.
-Zen, and S.
-Sorella, TurboRVB: A many-body toolkit for ab initio electronic simulations by quantum Monte Carlo, The Journal of Chemical Physics 152, 204121 (2020).
+(2020) Kousuke Nakano, Claudio Attaccalite, Matteo Barborini, Luca Capriotti, Michele Casula, Emanuele Coccia, Mario Dagrada, Claudio Genovese, Ye Luo, Guglielmo Mazzola, Andrea Zen, and Sandro Sorella, “TurboRVB: A many-body toolkit for ab initio electronic simulations by quantum Monte Carlo,” The Journal of Chemical Physics 152, 204121 (2020).
+- Xie et al. (2023a) Hao Xie, Zi-Hang Li, Han Wang, Linfeng Zhang, and Lei Wang, “Deep variational free energy approach to dense hydrogen,” Phys. Rev. Lett. 131, 126501 (2023a).
 - Morales et al.
-[2010b] M.
-A.
-Morales, C.
-Pierleoni, E.
-Schwegler, and D.
+(2010b) Miguel A.
+Morales, Carlo Pierleoni, Eric Schwegler, and D.
 M.
-Ceperley, Evidence for a first-order liquid-liquid transition in high-pressure hydrogen from ab initio simulations, Proceedings of the National Academy of Sciences 107, 12799 (2010b).
+Ceperley, “Evidence for a first-order liquid-liquid transition in high-pressure hydrogen from ab initio simulations,” Proceedings of the National Academy of Sciences 107, 12799–12803 (2010b).
 - Niu et al.
-[2023] H.
-Niu, Y.
-Yang, S.
-Jensen, M.
-Holzmann, C.
-Pierleoni, and D.
-M.
-Ceperley, Stable solid molecular hydrogen above 900 k from a machine-learned potential trained with diffusion quantum monte carlo, Phys.
+(2023) Hongwei Niu, Yubo Yang, Scott Jensen, Markus Holzmann, Carlo Pierleoni, and David M.
+Ceperley, “Stable solid molecular hydrogen above 900 k from a machine-learned potential trained with diffusion quantum monte carlo,” Phys.
 Rev.
@@
 - Wang et al.
-[2023] X.
-Wang, Z.
-Wang, P.
-Gao, C.
-Zhang, J.
-Lv, H.
-Wang, H.
-Liu, Y.
-Wang, and Y.
-Ma, Data-driven prediction of complex crystal structures of dense lithium, Nature Communications 14, 2924 (2023).
-- Zwanzig [1954] R. W. Zwanzig, High‐Temperature Equation of State by a Perturbation Method. I. Nonpolar Gases, The Journal of Chemical Physics 22, 1420 (1954).
-- Bennett [1976] C. H. Bennett, Efficient estimation of free energy differences from monte carlo data, Journal of Computational Physics 22, 245 (1976).
+(2023) Xiaoyang Wang, Zhenyu Wang, Pengyue Gao, Chengqian Zhang, Jian Lv, Han Wang, Haifeng Liu, Yanchao Wang, and Yanming Ma, “Data-driven prediction of complex crystal structures of dense lithium,” Nature Communications 14, 2924 (2023).
+- Zwanzig (1954) Robert W. Zwanzig, “High‐Temperature Equation of State by a Perturbation Method. I. Nonpolar Gases,” The Journal of Chemical Physics 22, 1420–1426 (1954).
+- Bennett (1976) Charles H Bennett, “Efficient estimation of free energy differences from monte carlo data,” Journal of Computational Physics 22, 245–268 (1976).
+- Wirnsberger et al.
+(2020) Peter Wirnsberger, Andrew J.
+Ballard, George Papamakarios, Stuart Abercrombie, Sébastien Racanière, Alexander Pritzel, Danilo Jimenez Rezende, and Charles Blundell, “Targeted free energy estimation via learned mappings,” The Journal of Chemical Physics 153, 144112 (2020).
+- Wirnsberger et al.
+(2022) Peter Wirnsberger, George Papamakarios, Borja Ibarz, Sébastien Racanière, Andrew J Ballard, Alexander Pritzel, and Charles Blundell, “Normalizing flows for atomic solids,” Machine Learning: Science and Technology 3, 025009 (2022).
+- Ahmad and Cai (2022) Rasool Ahmad and Wei Cai, “Free energy calculation of crystalline solids using normalizing flows,” Modelling and Simulation in Materials Science and Engineering 30, 065007 (2022).
+- Molina-Taborda et al.
+(2024) Ana Molina-Taborda, Pilar Cossio, Olga Lopez-Acevedo, and Marylou Gabrié, “Active Learning of Boltzmann Samplers and Potential Energies with Quantum Mechanical Accuracy,” Journal of Chemical Theory and Computation 20, 8833–8843 (2024).
+- Li and Wang (2018) Shuo-Hui Li and Lei Wang, “Neural network renormalization group,” Phys. Rev. Lett. 121, 260601 (2018).
+- Wu et al. (2019) Dian Wu, Lei Wang, and Pan Zhang, “Solving statistical mechanics using variational autoregressive networks,” Phys. Rev. Lett. 122, 080602 (2019).
+- Xie et al.
+(2022) Hao Xie, Linfeng Zhang, and Lei Wang, “Ab-initio study of interacting fermions at finite temperature with neural canonical transformation,” Journal of Machine Learning 1, 38–59 (2022).
+- Xie et al. (2023b) Hao Xie, Linfeng Zhang, and Lei Wang, “ of two-dimensional electron gas: A neural canonical transformation study,” SciPost Phys. 14, 154 (2023b).
 - Chen et al.
-[2018] R.
-T.
+(2018) Ricky T.
 Q.
-Chen, Y.
-Rubanova, J.
-Bettencourt, and D.
-Duvenaud, Neural ordinary differential equations, in Proceedings of the 32nd International Conference on Neural Information Processing Systems, NIPS’18 (Curran Associates Inc., Red Hook, NY, USA, 2018) p.
+Chen, Yulia Rubanova, Jesse Bettencourt, and David Duvenaud, “Neural ordinary differential equations,” in Proceedings of the 32nd International Conference on Neural Information Processing Systems, NIPS’18 (Curran Associates Inc., Red Hook, NY, USA, 2018) p.
 6572–6583.
-- Zhang et al. [2018] L. Zhang, W. E, and L. Wang, Monge-ampère flow for generative modeling (2018), arXiv:1809.10188 [cs.LG] .
+- Zhang et al. (2018) Linfeng Zhang, Weinan E, and Lei Wang, “Monge-ampère flow for generative modeling,” (2018), arXiv:1809.10188 [cs.LG] .
 - Vaswani et al.
-[2017] A.
-Vaswani, N.
-Shazeer, N.
-Parmar, J.
-Uszkoreit, L.
-Jones, A.
-N.
-Gomez, L.
-u.
-Kaiser, and I.
-Polosukhin, Attention is all you need, in Advances in Neural Information Processing Systems, Vol.
+(2017) Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones, Aidan N Gomez, Ł ukasz Kaiser, and Illia Polosukhin, “Attention is all you need,” in Advances in Neural Information Processing Systems, Vol.
 30, edited by I.
 Guyon, U.
-V.
-Luxburg, S.
+Von Luxburg, S.
 Bengio, H.
@@
 Garnett (Curran Associates, Inc., 2017).
-- Köhler et al. [2019] J. Köhler, L. Klein, and F. Noé, Equivariant flows: sampling configurations for multi-body systems with symmetric energies (2019), arXiv:1910.00753 [stat.ML] .
+- Köhler et al. (2019) Jonas Köhler, Leon Klein, and Frank Noé, “Equivariant flows: sampling configurations for multi-body systems with symmetric energies,” (2019), arXiv:1910.00753 [stat.ML] .
 - Satorras et al.
-[2021] V.
-G.
-Satorras, E.
-Hoogeboom, and M.
-Welling, E(n) equivariant graph neural networks, in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol.
-139, edited by M.
-Meila and T.
-Zhang (PMLR, 2021) pp.
+(2021) Víctor Garcia Satorras, Emiel Hoogeboom, and Max Welling, “E(n) equivariant graph neural networks,” in Proceedings of the 38th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol.
+139, edited by Marina Meila and Tong Zhang (PMLR, 2021) pp.
 9323–9332.
 - Taddei et al.
-[2015] M.
-Taddei, M.
-Ruggeri, S.
-Moroni, and M.
-Holzmann, Iterative backflow renormalization procedure for many-body ground-state wave functions of strongly interacting normal fermi liquids, Phys.
+(2015) Michele Taddei, Michele Ruggeri, Saverio Moroni, and Markus Holzmann, “Iterative backflow renormalization procedure for many-body ground-state wave functions of strongly interacting normal fermi liquids,” Phys.
 Rev.
@@
 - Pfau et al.
-[2020] D.
-Pfau, J.
-S.
-Spencer, A.
-G.
+(2020) David Pfau, James S.
+Spencer, Alexander G.
 D.
@@
 C.
-Foulkes, Ab initio solution of the many-electron schrödinger equation with deep neural networks, Phys.
+Foulkes, “Ab initio solution of the many-electron schrödinger equation with deep neural networks,” Phys.
 Rev.
@@
 2, 033429 (2020).
-- Pescia et al. [2022] G. Pescia, J. Han, A. Lovato, J. Lu, and G. Carleo, Neural-network quantum states for periodic systems in continuous space, Phys. Rev. Res. 4, 023138 (2022).
+- Pescia et al.
+(2022) Gabriel Pescia, Jiequn Han, Alessandro Lovato, Jianfeng Lu, and Giuseppe Carleo, “Neural-network quantum states for periodic systems in continuous space,” Phys.
+Rev.
+Res.
+4, 023138 (2022).
 - Wilson et al.
-[2023] M.
-Wilson, S.
-Moroni, M.
-Holzmann, N.
-Gao, F.
-Wudarski, T.
-Vegge, and A.
-Bhowmik, Neural network ansatz for periodic wave functions and the homogeneous electron gas, Phys.
+(2023) Max Wilson, Saverio Moroni, Markus Holzmann, Nicholas Gao, Filip Wudarski, Tejs Vegge, and Arghya Bhowmik, “Neural network ansatz for periodic wave functions and the homogeneous electron gas,” Phys.
 Rev.
@@
 - Cassella et al.
-[2023] G.
-Cassella, H.
-Sutterud, S.
-Azadi, N.
+(2023) Gino Cassella, Halvard Sutterud, Sam Azadi, N.
 D.
-Drummond, D.
-Pfau, J.
-S.
+Drummond, David Pfau, James S.
 Spencer, and W.
@@
 C.
-Foulkes, Discovering quantum phase transitions with fermionic neural networks, Phys.
+Foulkes, “Discovering quantum phase transitions with fermionic neural networks,” Phys.
 Rev.
@@
 130, 036401 (2023).
-- Pescia et al. [2024] G. Pescia, J. Nys, J. Kim, A. Lovato, and G. Carleo, Message-passing neural quantum states for the homogeneous electron gas, Phys. Rev. B 110, 035108 (2024).
+- Pescia et al.
+(2024) Gabriel Pescia, Jannes Nys, Jane Kim, Alessandro Lovato, and Giuseppe Carleo, “Message-passing neural quantum states for the homogeneous electron gas,” Phys.
+Rev.
+B 110, 035108 (2024).
 - Giannozzi et al.
-[2009] P.
-Giannozzi, S.
-Baroni, N.
-Bonini, M.
-Calandra, R.
-Car, C.
-Cavazzoni, D.
-Ceresoli, G.
-L.
-Chiarotti, M.
-Cococcioni, I.
-Dabo, A.
-Dal Corso, S.
-de Gironcoli, S.
-Fabris, G.
-Fratesi, R.
-Gebauer, U.
-Gerstmann, C.
-Gougoussis, A.
-Kokalj, M.
-Lazzeri, L.
-Martin-Samos, N.
-Marzari, F.
-Mauri, R.
-Mazzarello, S.
-Paolini, A.
-Pasquarello, L.
-Paulatto, C.
-Sbraccia, S.
-Scandolo, G.
-Sclauzero, A.
-P.
-Seitsonen, A.
-Smogunov, P.
-Umari, and R.
-M.
-Wentzcovitch, Quantum espresso: a modular and open-source software project for quantum simulations of materials, Journal of Physics: Condensed Matter 21, 395502 (19pp) (2009).
-- Blöchl [1994] P. E. Blöchl, Projector augmented-wave method, Phys. Rev. B 50, 17953 (1994).
-- Monkhorst and Pack [1976] H. J. Monkhorst and J. D. Pack, Special points for brillouin-zone integrations, Phys. Rev. B 13, 5188 (1976).
-- Bussi et al. [2007] G. Bussi, D. Donadio, and M. Parrinello, Canonical sampling through velocity rescaling, The Journal of Chemical Physics 126, 014101 (2007).
-- Chen and Lipman [2024] R. T. Q. Chen and Y. Lipman, Flow matching on general geometries, in The Twelfth International Conference on Learning Representations (2024).
-- Klein et al. [2023] L. Klein, A. Krämer, and F. Noe, Equivariant flow matching, in Thirty-seventh Conference on Neural Information Processing Systems (2023).
-- Kuhn [1955] H. W. Kuhn, The hungarian method for the assignment problem, Naval Research Logistics Quarterly 2, 83 (1955).
-- Militzer [2013] B. Militzer, Equation of state calculations of hydrogen-helium mixtures in solar and extrasolar giant planets, Phys. Rev. B 87, 014202 (2013).
-- Becker et al. [2013] A. Becker, N. Nettelmann, B. Holst, and R. Redmer, Isentropic compression of hydrogen: Probing conditions deep in planetary interiors, Phys. Rev. B 88, 045122 (2013).
+(2009) Paolo Giannozzi, Stefano Baroni, Nicola Bonini, Matteo Calandra, Roberto Car, Carlo Cavazzoni, Davide Ceresoli, Guido L Chiarotti, Matteo Cococcioni, Ismaila Dabo, Andrea Dal Corso, Stefano de Gironcoli, Stefano Fabris, Guido Fratesi, Ralph Gebauer, Uwe Gerstmann, Christos Gougoussis, Anton Kokalj, Michele Lazzeri, Layla Martin-Samos, Nicola Marzari, Francesco Mauri, Riccardo Mazzarello, Stefano Paolini, Alfredo Pasquarello, Lorenzo Paulatto, Carlo Sbraccia, Sandro Scandolo, Gabriele Sclauzero, Ari P Seitsonen, Alexander Smogunov, Paolo Umari, and Renata M Wentzcovitch, “Quantum espresso: a modular and open-source software project for quantum simulations of materials,” Journal of Physics: Condensed Matter 21, 395502 (19pp) (2009).
+- Blöchl (1994) P. E. Blöchl, “Projector augmented-wave method,” Phys. Rev. B 50, 17953–17979 (1994).
+- Monkhorst and Pack (1976) Hendrik J. Monkhorst and James D. Pack, “Special points for brillouin-zone integrations,” Phys. Rev. B 13, 5188–5192 (1976).
+- Bussi et al. (2007) Giovanni Bussi, Davide Donadio, and Michele Parrinello, “Canonical sampling through velocity rescaling,” The Journal of Chemical Physics 126, 014101 (2007).
+- Chen and Lipman (2024) Ricky T. Q. Chen and Yaron Lipman, “Flow matching on general geometries,” in The Twelfth International Conference on Learning Representations (2024).
+- Klein et al. (2023) Leon Klein, Andreas Krämer, and Frank Noe, “Equivariant flow matching,” in Thirty-seventh Conference on Neural Information Processing Systems (2023).
+- Kuhn (1955) H. W. Kuhn, “The hungarian method for the assignment problem,” Naval Research Logistics Quarterly 2, 83–97 (1955).
+- Saumon and Chabrier (1991) Didier Saumon and Gilles Chabrier, “Fluid hydrogen at high density: Pressure dissociation,” Phys. Rev. A 44, 5122–5141 (1991).
+- Militzer (2013) Burkhard Militzer, “Equation of state calculations of hydrogen-helium mixtures in solar and extrasolar giant planets,” Phys. Rev. B 87, 014202 (2013).
+- Becker et al.
+(2013) Andreas Becker, Nadine Nettelmann, Bastian Holst, and Ronald Redmer, “Isentropic compression of hydrogen: Probing conditions deep in planetary interiors,” Phys.
+Rev.
+B 88, 045122 (2013).
 - Morales et al.
-[2013] M.
-A.
-Morales, J.
-M.
-McMahon, C.
-Pierleoni, and D.
-M.
-Ceperley, Nuclear quantum effects and nonlocal exchange-correlation functionals applied to liquid hydrogen at high pressure, Phys.
+(2013) Miguel A.
+Morales, Jeffrey M.
+McMahon, Carlo Pierleoni, and David M.
+Ceperley, “Nuclear quantum effects and nonlocal exchange-correlation functionals applied to liquid hydrogen at high pressure,” Phys.
 Rev.
@@
 110, 065702 (2013).
-- Chabrier and Potekhin [1998] G. Chabrier and A. Y. Potekhin, Equation of state of fully ionized electron-ion plasmas, Phys. Rev. E 58, 4941 (1998).
-- Militzer and Ceperley [2000] B. Militzer and D. M. Ceperley, Path integral monte carlo calculation of the deuterium hugoniot, Phys. Rev. Lett. 85, 1890 (2000).
-- Militzer and Ceperley [2001] B. Militzer and D. M. Ceperley, Path integral monte carlo simulation of the low-density hydrogen plasma, Phys. Rev. E 63, 066404 (2001).
+- Chabrier and Potekhin (1998) Gilles Chabrier and Alexander Y. Potekhin, “Equation of state of fully ionized electron-ion plasmas,” Phys. Rev. E 58, 4941–4949 (1998).
+- Militzer and Ceperley (2000) B. Militzer and D. M. Ceperley, “Path integral monte carlo calculation of the deuterium hugoniot,” Phys. Rev. Lett. 85, 1890–1893 (2000).
+- Militzer and Ceperley (2001) B. Militzer and D. M. Ceperley, “Path integral monte carlo simulation of the low-density hydrogen plasma,” Phys. Rev. E 63, 066404 (2001).
 Supplemental Material
@@
 In this work, we use the flow matching method to benchmark the calculation of free energy difference (scaled by temperature) between the two states, , against the conventional thermodynamic integration (TI) approach.
-Below we will discuss various aspect of the new method, including the underlying theoretical framework, training objectives, implementation details, and some illustrative examples of the results.
+Below we will discuss various aspects of the new method, including the underlying theoretical framework, training objectives, implementation details, and some illustrative examples of the results.
 Appendix A Targeted free energy perturbation
-In contrast to TI, flow matching is built on the framework of targeted free energy perturbation (TFEP) [28, 29], as already mentioned in the main text.
-Compared to the original free energy perturbation based on importance sampling [58], TFEP further leverages an invertible map in the configuration space to significantly increase state overlap and thereby reduces the variance of relevant estimators.
+In contrast to TI, flow matching is built on the framework of targeted free energy perturbation (TFEP) Jarzynski (2002); Hahn and Then (2009), as already mentioned in the main text.
+Compared to the original free energy perturbation based on importance sampling Zwanzig (1954), TFEP further leverages an invertible map in the configuration space to significantly increase state overlap and thereby reduces the variance of relevant estimators.
 More specifically, consider a map that transforms coordinate samples of the state 0 into a new state , whose probability density function is given by the change-of-variables formula:
@@
 are known as the forward and reverse work, respectively. These results are precisely the Eqs. (2) and (3) in the main text.
-Note that due to limited number of samples in practice, the two estimators in Eq.
-(S4) are actually biased, which can be readily demonstrated using Jensen’s inequality [29].
+Note that due to the limited number of samples in practice, the two estimators in Eq.
+(S4) are actually biased, which can be readily demonstrated using Jensen’s inequality Hahn and Then (2009).
 Specifically, the estimator involving the forward (reverse) work turns out to be an upper (lower) bound of the true free energy difference value.
@@
 In this work, we found these two bounds are generally quite tight; see the sections below for some illustrative examples.
-As a result, we choose to simply perform an arithmetic mean of the two bounds to generate the final data, although some more sophisticated and potentially superior approaches may also be considered [59].
-Another observation is that the estimation of absolute entropy appears to be more straightforward within the TFEP framework compared to TI: once one of the two states is chosen to have tractable entropy, the practical implementation remains almost unchanged.
+As a result, we choose to simply perform an arithmetic mean of the two bounds to generate the final free energy and entropy data, although some more sophisticated and potentially superior approaches may also be considered Bennett (1976).
+Another observation is that the estimation of absolute entropy appears to be more straightforward within the TFEP framework compared to TI: once one of the two states is chosen to have known entropy, the practical implementation remains almost unchanged.
 Appendix B Flow matching
@@
 The recent progress in the field of deep generative modeling has largely changed the situation.
-Specifically, the desired invertible map can be efficiently represented by a class of generative models called normalizing flow [30], making TFEP reemerge as a promising tool for free energy calculation.
-In previous attempts [31, 32, 33, 34], the training objectives have been exclusively based on variants of the Kullback-Leibler (KL) divergence, which provide direct access to the free energy difference as a variational upper or lower bound.
-This variational training scheme has been successfully applied to a broad range of physical problems, including lattice models [35, 36], atomic solids [32, 33], quantum dots [37], uniform electron gases [38], and also dense hydrogen [39].
+Specifically, the desired invertible map can be efficiently represented by a class of generative models called normalizing flow Papamakarios et al.
+(2021), making TFEP reemerge as a promising tool for free energy calculation.
+In previous attempts Wirnsberger et al.
+(2020, 2022); Ahmad and Cai (2022); Molina-Taborda et al.
+(2024), the training objectives have been exclusively based on variants of the Kullback-Leibler (KL) divergence, which provide direct access to the free energy difference as a variational upper or lower bound.
+This variational training scheme has been successfully applied to a broad range of physical problems, including lattice models Li and Wang (2018); Wu et al.
+(2019), atomic solids Wirnsberger et al.
+(2022); Ahmad and Cai (2022), quantum dots Xie et al.
+(2022), uniform electron gases Xie et al.
+(2023b), and also dense hydrogen Xie et al.
+(2023a).
 However, it still has some drawbacks for the present purpose.
 First, the training process typically requires a large number of energy calculations of the underlying system, which would be time-consuming for ab initio methods like DFT.
-As a result, many works have used empirical [31, 32, 33] or machine-learned potentials [34], which would inevitably introduce new source of errors.
-Second, symmetry constraints often play an essential role in the efficient training of flow model, especially for systems like pure hydrogen.
-Although several neural network designs have been proposed to meet various symmetry requirements [31, 32], they are still not satisfactory enough regarding a proper balance between the expressive power and computational cost.
-The continuous normalizing flow (CNF) [60, 61] is arguably the most convenient and flexible candidate now to incorporate symmetries, but its variational training requires expensive numerical integrations of ordinary differential equation (ODE) and scales unfavorably for large system sizes.
-The recently developed flow matching method [23, 24, 25] has largely overcome these difficulties.
+As a result, many works have used empirical Wirnsberger et al.
+(2020, 2022); Ahmad and Cai (2022) or machine-learned potentials Molina-Taborda et al.
+(2024), which would inevitably introduce new source of errors.
+Second, symmetry constraints often play an essential role in the efficient training of flow model, especially for homogeneous systems like pure hydrogen.
+Although several neural network designs have been proposed to meet various symmetry requirements Wirnsberger et al.
+(2020, 2022), they are still not satisfactory enough regarding a proper balance between the expressive power and computational cost.
+The continuous normalizing flow (CNF) Chen et al.
+(2018); Zhang et al.
+(2018) is arguably the most convenient and flexible candidate now to incorporate symmetries, but its variational training requires expensive numerical integrations of ordinary differential equation (ODE) and scales unfavorably for large system sizes.
+The recently developed flow matching method Albergo and Vanden-Eijnden (2023); Lipman et al.
+(2023); Liu et al.
+(2023) has largely overcome these difficulties.
 More specifically, one can smoothly connect the two end probability distributions by introducing an interpolant , where is a continuous “time” parameter, such as
 The essence is that the velocity field underlying such an interpolated “probability path” can then be obtained by minimization of the following objective:
-Notice this expression requires only samples from the two end distributions; there are not any on-the-fly evaluation of configurational energy or ODE integration during the training stage.
+Notice this expression requires only samples from the two end distributions; there are not any on-the-fly evaluations of configurational energy or ODE integration during the training stage.
 On the other hand, these types of costly computations are indeed present when estimating the free energy difference after the training is finished, as shown in Eqs.
 (S4) and (S5).
-In particular, the transformed coordinates and corresponding log-Jacobian determinant should be evaluated by jointly integrating the following set of ODEs [60, 61]:
+In particular, the transformed coordinates and corresponding log-Jacobian determinant should be evaluated by jointly integrating the following set of ODEs Chen et al.
+(2018); Zhang et al.
+(2018):
 However, it should be noted that such computations are performed only once and thus still practically affordable.
-As mentioned above, compared to previous network designs for discrete flow model [31, 32], the incorporation of various symmetries can also be more convenient for continuous normalizing flow.
+As mentioned above, compared to previous network designs for discrete flow model Wirnsberger et al.
+(2020, 2022), the incorporation of various symmetries can also be more convenient for continuous normalizing flow.
 Specifically, the invariance of probability distribution can simply be achieved by requiring the underlying velocity field to be equivariant under the desired symmetry operation :
-To do this in practice, we can leverage many recent advances in various areas, such as natural language processing [62], molecular simulation [63, 31], graph neural networks [64], and quantum many body computations [65, 66, 38, 67, 39, 68, 69, 70].
+To do this in practice, we can leverage many recent advances in various areas, such as natural language processing Vaswani et al.
+(2017), molecular simulation Köhler et al.
+(2019); Wirnsberger et al.
+(2020), graph neural networks Satorras et al.
+(2021), and quantum many body computations Taddei et al.
+(2015); Pfau et al.
+(2020); Xie et al.
+(2023b); Pescia et al.
+(2022); Xie et al.
+(2023a); Wilson et al.
+(2023); Cassella et al.
+(2023); Pescia et al.
+(2024).
 Appendix C Implementation details
-The MD generation of coordinate samples used for training Eq.
+The MD generation of coordinate samples used for optimizing Eq.
 (S7) and the energy computation appearing in the free energy estimators Eq.
-(S4) are both performed in the DFT level using the QUANTUM ESPRESSO code [71].
-We use the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) [15] and the pseudopotential of projector-augmented wave [72] type.
-We choose the system size , and the -space integration is performed on a Monkhorst-Pack grid [73].
+(S4) are both performed in the DFT level using the QUANTUM ESPRESSO code Giannozzi et al.
+(2009).
+We use the exchange-correlation functional of Perdew, Burke and Ernzerhof (PBE) Perdew et al.
+(1996) and the pseudopotential of projector-augmented wave Blöchl (1994) type.
+We choose the system size , and the -space integration is performed on a Monkhorst-Pack grid Monkhorst and Pack (1976).
 The energy cutoff for the plane wave basis set and charge densities are and Ry, respectively.
-These parameters have been checked to yield well converged energetics.
-For MD simulations, the ion temperature is controlled using stochastic-velocity rescaling [74] and the time step is set to 12 a.u..
-The network we use for the velocity field is adapted from FermiNet [66], which is permutation and translation equivariant by construction.
-We also modify the pair distance features to comply with the periodic nature of the simulation box [38, 67, 68, 69].
+Note the parameters above have been checked to yield well converged pressures and energies.
+For MD simulations, the ion temperature is controlled using stochastic-velocity rescaling Bussi et al.
+(2007) and the time step is set to 12 a.u..
+The network we use for the velocity field is adapted from FermiNet Pfau et al.
+(2020), which is permutation and translation equivariant by construction.
+We also modify the pair distance features to comply with the periodic nature of the simulation box Xie et al.
+(2023b); Pescia et al.
+(2022); Wilson et al.
+(2023); Cassella et al.
+(2023).
 We choose a simple linear interpolating function
@@
 This setting turns out to be equivalent to the Riemannian flow matching on a torus as proposed in Ref.
-[75].
+Chen and Lipman (2024).
 To allow for more efficient use of the training data, we choose to further incorporate permutation symmetry explicitly into the flow matching objective Eq.
-(S7) itself [76].
-In practice, this can be achieved by finding the permutation with minimal distance for each pair of samples using the Hungarian algorithm [77].
+(S7) itself Klein et al.
+(2023).
+In practice, this can be achieved by finding the permutation with minimal distance for each pair of samples using the Hungarian algorithm Kuhn (1955).
 Appendix D Illustrative examples of the results
@@
 This is a direct consequence of the fluctuation theorem as derived in Ref.
-[29].
+Hahn and Then (2009).
 We also use the same workflow outlined above to calculate the absolute free energy and entropy of the reference state at and , as mentioned in the main text.
@@
 Nonetheless, the absolute entropy of the reference state can still be estimated with a relative error of , which is comparable to the results reported by Morales et al.
-[19] using Hamiltonian thermodynamic integration and already satisfactory for the present purpose (see Fig.
+Morales et al.
+(2010a) using Hamiltonian thermodynamic integration and already satisfactory for the present purpose (see Fig.
 1(a) in the main text).
@@
 S5 also shows the entropy reported by Miguel et al.
-[4] (dashed lines).
+Miguel, Y.
+et al.
+(2016) (dashed lines).
 According to the authors, they processed the original REOS3 data by integrating along isotherm first, hence in accordance with the red solid lines of our result.
 However, these two sets of entropies clearly differ by an offset in the ab initio region of the main panel.
-The remedy of such an apparent disagreement turns out to be surprisingly easy: notice all the entropy isochores already align well, we just need to locate the reference point on another “correct” isotherm, say, 15000K.
+The remedy of this apparent disagreement turns out to be surprisingly easy: notice all the entropy isochores already align well, we just need to locate the reference point on another “correct” isotherm, say, 15000K.
 After doing this, the red solid lines then coincide perfectly with the dashed lines from Ref.
-[4], as shown in Figure S6.
+Miguel, Y.
+et al.
+(2016), as shown in Figure S6.
 Our reproduction of the entropy results by Miguel et al.
-[4] based on a specific integration path and reference point, as done above, demonstrates how fragile the “global” TI approach can be for an EOS with significant thermodynamic inconsistencies.
+Miguel, Y.
+et al.
+(2016) based on an ad hoc choice of the integration path and reference point, as done above, demonstrates how fragile the “global TI” approach can be for an EOS with significant thermodynamic inconsistencies.
 Notice the solid and dashed lines in Fig.
@@
 To construct an EOS that is thermodynamically consistent across several distinct theories, such as ab initio DFT and SCvH, one needs to first calculate the free energy and entropy of each region separately using the standard TI approach.
-We then perform a two-dimensional spline interpolation over each set of free energy data, following a similar approach to Militzer and Hubbard [12].
+We then perform a two-dimensional spline interpolation over each set of free energy data, following a similar approach to Militzer and Hubbard Militzer and Hubbard (2013).
 This in principle allows us to reproduce the original pressure and energy data by performing partial derivatives.
@@
 This indicates that we are allowed to globally shift the DFT energies as we want, which, of course, leaves the entropy (and any other thermodynamic properties) unchanged.
-Second, we have surprisingly found the SCvH entropies are about (or ) higher than well-established experimental data at room conditions [46, 47], as mentioned in the main text.
-As shown in Figure S7, this discrepancy between the two sets of data is fairly uniform across a broad region of (low) temperatures and densities, including the standard condition (1 bar, 298.15K) that has been marked explicitly.
-To get a feeling of the impact such an entropy error can have on the interpolation, Figure S8 shows the resulting EOS using the original (but incorrect) SCvH entropy data.
-Note that although we have selected an DFT energy offset to connect the free energy isotherms on both sides as smoothly as possible (see panel (c)), there still remains a noticeable mismatch.
+Second, note that in the DFT-MD region, we have implicitly neglected the entropy contribution from the proton spin degrees of freedom, as they are irrelevant for essentially classical protons at high temperatures.
+However, this constant plays an important role when trying to connect with SCvH, which has taken account of accurate energy levels arising from both proton spin isomers (i.e., para- and ortho-hydrogen) when constructing the partition function of molecular hydrogen gas at low densities Saumon and Chabrier (1991).
+As a result, the constant must be subtracted from the original SCvH data to ensure a truly fair matching of the absolute entropies on both sides.
+This is precisely what we have done to produce the final EOS shown in Fig.
+3 of the main text.
+To get a feeling on how good is the interpolation quality, it is instructive and interesting to see how bad it will become if we just mindlessly adopt the original SCvH entropy data without subtracting the constant mentioned above.
+The resulting EOS based on the same interpolation procedure is shown in Figure S7.
+Notice that although we have selected a DFT energy offset to connect the free energy isotherms on both sides as smoothly as possible (see panel (c)), there still remains a noticeable mismatch.
+In fact, according to the basic relation , the free energy isotherms of SCvH are apparently wider than the ab initio region due to too large entropy on the SCvH side.
 On the other hand, Fig.
-S8(d) shows the interpolated pressure is also apparently problematic.
-In particular, the wild wiggling behavior at the intermediate densities can even cause violation of the basic mechanical stability constraint [21] and hence definitely unphysical.
-Notice that in Fig.
-S8(c), the free energy isotherms of SCvH appear to be wider than the ab initio region.
-Recalling the basic relation , the problematic interpolation can then be remedied by globally reducing the SCvH entropy to match the correct experimental value.
-This leads to the final EOS shown in Fig.
-3 of the main text.
-Note we have also added an extra entropy offset of on the DFT side to make the interpolation look slightly better.
-Nevertheless, this residual error is much smaller than the SCvH entropy error discussed above and does not affect our essential arguments on the accuracy and reliability of the new EOS construction approach.
-In principle, the remaining small error near the interpolation boundaries can be eliminated by using more accurate chemical models at low densities or electron structure methods (such as QMC) at high densities.
+S7(d) shows the interpolated pressure is also problematic.
+In particular, the wild wiggling behavior at the intermediate densities can even cause violation of the basic mechanical stability constraint Saumon et al.
+(1995) and hence definitely unphysical.
+The observations above indicate that the interpolation quality serves as a very sensitive and stringent “detector” of the accuracy of both theories.
+In particular, we have been forced to identify, confirm and “fix” the constant shift in original SCvH entropy data, even if we were totally unaware of this subtlety in the early stages of this work.
+This is a very compelling demonstration of the power of our new EOS construction protocol.
+In particular, such an iteratively “self-correcting” mechanism possessed by the protocol is exactly what we need to ensure steady convergence towards a conclusive hydrogen EOS for planetary modeling.
+Although we have achieved a fairly excellent matching between SCvH and DFT-PBE-MD, there still remains some small errors.
+In particular, note we actually have also added an extra (and very small) entropy offset of on the DFT side to make the interpolation shown in Fig.
+3 of the main text look slightly better.
+In principle, such remaining errors can be eliminated by using more accurate chemical models at low densities or electron structure methods (such as QMC) at high densities.
 Besides, the nuclear quantum effect may also have a slight impact on the ab initio entropy at low temperatures, which has been ignored in this work.
-This can be remedied either by adding some semi-empirical corrections [78, 79] or resorting to more accurate (and also more costly) path integral MD simulations [80].
+This can be remedied either by adding some semi-empirical corrections Militzer (2013); Becker et al.
+(2013) or resorting to more accurate (and also more costly) path integral MD simulations Morales et al.
+(2013).
 Appendix G More details of comparison between different Jupiter’s adiabats
@@
 Compared to Fig.
-4 of the main text, there is additionally the adiabat derived from the CMS19 EOS [11] (yellow line), which is clearly smoother in the ab initio region than both our result (red dashed line) and MH13 (blue line).
+4 of the main text, there is additionally the adiabat derived from the CMS19 EOS Chabrier et al.
+(2019) (yellow line), which is clearly smoother in the ab initio region than both our result (red dashed line) and MH13 (blue line).
 This difference originates from the significant, non-monotonic slope change of our entropy isotherms, as shown in Fig.
@@
 In Fig.
-S9, we also directly illustrate the entropy of our EOS by colored regions.
+S8, we also directly illustrate the entropy of our EOS by colored regions.
 The change of slope is again clearly evident, indicating that our EOS better captures the pressure-induced molecular dissociation than CMS19.
 By definition, our calculated adiabat (red dashed line) follows the shape of a certain “stripe” of constant entropy, as clearly shown in Fig.
-S9.
+S8.
 Note that the slightly imperfect behavior of our adiabat between 1 and 10GPa corresponds to the interpolation region of our EOS; see the previous section for more discussions.
-However, it is worth emphasizing that this remaining error will not affect the accuracy and reliability of our adiabat at higher pressures, where the disagreements between various adiabats make the biggest difference on resulting planetary model predictions [4, 7].
+However, it is worth emphasizing that this remaining error will not affect the accuracy and reliability of our adiabat at higher pressures, where the disagreements between various adiabats make the biggest difference on resulting planetary model predictions Miguel, Y.
+et al.
+(2016); Howard, S.
+et al.
+(2023).
 In Fig.
-S9, our adiabat is confined to the ab initio region explored in this work, specifically up to K and .
-It is in principle just a matter of additional DFT-MD simulations or interpolations with other theories [21, 81, 82, 83] to extend the hydrogen EOS to higher temperatures and densities.
+S8, our adiabat is confined to the ab initio region explored in this work, specifically up to K and .
+It is in principle just a matter of additional DFT-MD simulations or interpolations with other theories Saumon et al.
+(1995); Chabrier and Potekhin (1998); Militzer and Ceperley (2000, 2001) to extend the hydrogen EOS to higher temperatures and densities.
 This will enable us to accurately study Jupiter’s deeper interior beyond the red dashed line in Fig.
-S9, or other astrophysical objects under even more extreme conditions, such as solar-type stars.
+S8, or other astrophysical objects under even more extreme conditions, such as solar-type stars.